### v2

# Causal Explanations for Image Classifiers

Abstract
Existing algorithms for explaining the output of image classifiers use different definitions of explanations and a variety of techniques to extract them. However, none of the existing tools use a principled approach based on formal definitions of causes and explanations for the explanation extraction.
In this paper we present a novel black-box approach to computing explanations grounded in the theory of actual causality. We prove relevant theoretical results and present an algorithm for computing approximate explanations based on these definitions. We prove termination of our algorithm and discuss its complexity and the amount of approximation compared to the precise definition.
We implemented the framework in a tool rex and we present experimental results and a comparison with state-of-the-art tools. We demonstrate that rex is the most efficient tool and produces the smallest explanations, in addition to outperforming other black-box tools on standard quality measures.
1 Introduction
(a) Ladybug
(b) Responsibility
(c) Heat map
(d) Explanation
[…]
Neural networks (NNs) are now a primary building block of many computer vision systems. NNs are complex non-linear functions with algorithmically generated (and not engineered) coefficients. In contrast to traditionally engineered image processing pipelines it is difficult to retrace how the pixel data are interpreted by the layers of the network. Moreover, in many application areas, in particular healthcare, the networks are proprietary, making the analysis of the internal layers impossible. This “black box” nature of neural networks creates demand for techniques that explain why a particular input yields the output that is observed without understanding the model’s parameters and their influence on the output.
An explanation of an output of an automated procedure is essential in many areas, including verification, planning, diagnosis and the like. A good explanation can increase a user’s confidence in the result. Explanations are also useful for determining whether there is a fault in the automated procedure: if the explanation does not make sense, it may indicate that the procedure is faulty. It is less clear how to define what a good explanation is. There have been a number of definitions of explanations over the years in various domains of computer science (???), philosophy (?) and statistics (?).
Black-box explanations for the results of image classifiers are typically based on or are given in the form of a ranking of the pixels, which is a numerical measure of importance: the higher the score, the more important the pixel is for the NN’s classification outcome. Often these rankings are presented in a form of a heat map, with higher scores corresponding to hotter areas. A user-friendly explanation can then be a subset of highest-ranked pixels that is sufficient for the original classification outcome.
This paper addresses the following research questions:
RQ1
What is a formal rigorous definition of explanation, suitable for image classification?
RQ2
Can we compute explanations based on the definition above without opening the black box? What is the complexity of such computation?
RQ3
Is there an efficiently computable approximation to explanations?
RQ4
What are suitable quality measures for explanations? A measure is suitable if it is meaningful with respect to quality and produces results that match our intuition.
RQ5
What is the quality of explanations computed by our algorithms compared to other black-box methods?
RQ6
Is there a trade-off between quality and compute cost of the explanations?
RQ7
Can black-box methods achieve the same quality of explanations as white-box methods?
[…]
We compare the experimental results of our implementation of the algorithm in a tool rex with those of other black-box tools on a number of benchmark datasets: ImageNet, VOC2012, and a set of partially occluded images. Our results show that rex is on par with the most efficient tools in terms of execution time, while producing the smallest explanations. We also analyse the explanations using the standard measures of insertion and deletion curves and the intersection with the background or the occlusion. The results show that rex outperforms the other tools on insertion curves and the overlap with irrelevant parts of the image. We discuss the logic behind deletion curves and demonstrate that low deletion is a measure of the quality of the model for particular types of images, rather than for the quality of an explanation. We also argue that explanations computed by rex are so small that any further reduction in their size is meaningless.
The rest of the paper is organized as follows. Section 2 presents an overview of the related work on black-box explainable AI. Section 3 gives the necessary background on actual causality. Section 4 provides the theoretical foundations of our approach to explainability. Section 5 is a detailed overview of our algorithm, with the evaluation results discussed in Section 6. The tool rex is open source and available at https://github.com/ReX-XAI/ReX. All benchmark sets are publicly available and referenced in Section 6.
2 Related Work
There is a large body of work on explaining the results of image classifiers. As our primary interest is in black-box approaches, we only give a brief overview of the white-box algorithms here. The existing approaches to post hoc explainability can be largely grouped into two categories: propagation and perturbation.
Propagation-based explanation methods are often regarded as more efficient. They back-propagate a model’s decision to the input layer to determine the weight of each input feature for the decision. Grad-CAM by (?) only needs one backward pass and propagates the class-specific gradient into the final convolutional layer of a DNN to coarsely highlight important regions of an input image. Guided Back-Propagation (gbp) by (?) computes the single and average partial derivatives of the output to attribute the prediction of a DNN. Integrated Gradients (ig) by (?) further uses two axioms called sensitivity and completeness for the problem of how to attribute the classification by a deep network to its input features. noisetunnel by (?) adds gaussian noise to the image before applying another XAI method. The noise serves to smooth any resultant saliency map, improving explanation quality. shap (SHapley Additive exPlanations) by (?) goes beyond propagation and identifies inputs that are similar to the input for which the output is to be explained, and ranks the features of the input according to their difference. A key advantage of shap is that it does not require counterfactuals. gradientshap is gradient-based method that approximates Shapley values via gradient integration and a noise tunnel. Both ig and gradientshap require access to model gradients while computing explanations. They form part of a family of whitebox explanation methods which differ from the Grad-CAM family in that they do not rely on examination of specified activation layers in the model. Other methods for approximating Shapley values are numerous, among which KernelShap (?) and DeepLiftShap (?) are probably the most used. KernelShap uses the lime framework to approximate Shapley values, whereas DeepListShap uses the DeepLift (?) method to the same end.
In contrast to propagation-based explanation methods, perturbation-based explanation approaches explore the input space directly in search for an explanation. The exploration/search often requires a large number of inference passes, which incurs significant computational cost when compared to propagation methods. There are many sampling methods, most of which are based on random search or heuristics. The chief advantage of perturbation-based methods is that they are model independent and do not require any access to model internals. Given a particular input, lime by (?) samples the the neighborhood of this input and creates a inherently explainable model to approximate the system’s local behavior; owing to the high computational cost of this approach, the ranking uses super-pixels via segmentation instead of individual pixels. (?) replace the natural distribution of the input by a user-defined distribution and then use the Shapley Value method to analyze combinations of input features and to rank their importance. (?) estimate the importance of input features by measuring the the flow of information between inputs and outputs. Both the Shapley Value and the information-theoretic approaches are computationally expensive. In rise by (?), the importance of a pixel is computed as the expectation over all local perturbations conditioned on the event that the pixel is observed. (?) have introduced the concept of “extreme perturbations” to improve the perturbation analysis in the extremal algorithm. More recently, spectrum-based fault localisation (SBFL) has been applied to explaining image classifiers. The technique has been implemented in the tool deepcover by (?).
The method and tool, rex, presented in this paper is a perturbation-based approach and addresses the limitations of existing black-box methods in two aspects. The feature masking in rex uses causal reasoning that provides a guarantee (subject to the assumption). Owing to its guided iterative refinement, rex is computationally efficient, while still producing explanations that are the state of the art.
3 Background on Actual Causality
In this section, we briefly review the definitions of causality and causal models introduced by (?) and relevant definitions of causes and explanations in image classification by (?). The reader is referred to (?) for further reading.
We assume that the world is described in terms of variables and their values. Some variables may have a causal influence on others. This influence is modeled by a set of structural equations. It is conceptually useful to split the variables into two sets: the exogenous variables, whose values are determined by factors outside the model, and the endogenous variables, whose values are ultimately determined by the exogenous variables. The structural equations describe how these values are determined.
Formally, a causal model $M$ is a pair $({\mathcal{}S},{\mathcal{}F})$, where ${\mathcal{}S}$ is a signature, which explicitly lists the endogenous and exogenous variables and characterizes their possible values, and ${\mathcal{}F}$ defines a set of (modifiable) structural equations, relating the values of the variables. A signature ${\mathcal{}S}$ is a tuple $({\mathcal{}U},{\mathcal{}V},{\mathcal{}R})$, where ${\mathcal{}U}$ is a set of exogenous variables, ${\mathcal{}V}$ is a set of endogenous variables, and ${\mathcal{}R}$ associates with every variable $Y\in{\mathcal{}U}\cup{\mathcal{}V}$ a nonempty set ${\mathcal{}R}(Y)$ of possible values for $Y$ (i.e., the set of values over which $Y$ ranges). For simplicity, we assume here that ${\mathcal{}V}$ is finite, as is ${\mathcal{}R}(Y)$ for every endogenous variable $Y\in{\mathcal{}V}$. The set ${\mathcal{}F}$ associates with each endogenous variable $X\in{\mathcal{}V}$ a function denoted $F_{X}$ (i.e., $F_{X}={\mathcal{}F}(X)$) such that $F_{X}:(\times_{U\in{\mathcal{}U}}{\mathcal{}R}(U))\times(\times_{Y\in{\mathcal{}V}-\{X\}}{\mathcal{}R}(Y))\rightarrow{\mathcal{}R}(X)$.
The structural equations define what happens in the presence of external interventions. Setting the value of some variable $X$ to $x$ in a causal model $M=({\mathcal{}S},{\mathcal{}F})$ results in a new causal model, denoted $M_{X\leftarrow x}$, which is identical to $M$, except that the equation for $X$ in ${\mathcal{}F}$ is replaced by $X=x$.
Probabilistic causal models are pairs $(M,\Pr)$, where $M$ is a causal model and $\Pr$ is a probability on the contexts. A causal model $M$ is recursive (or acyclic) if its causal graph is acyclic. If $M$ is an acyclic causal model, then given a context, that is, a setting $\vec{u}$ for the exogenous variables in ${\mathcal{}U}$, the values of all the other variables are determined. In this paper we restrict the discussion to recursive models.
We call a pair $(M,\vec{u})$ consisting of a causal model $M$ and a context $\vec{u}$ a (causal) setting. A causal formula $\psi$ is true or false in a setting. We write $(M,\vec{u})\models\psi$ if the causal formula $\psi$ is true in the setting $(M,\vec{u})$. The $\models$ relation is defined inductively. $(M,\vec{u})\models X=x$ if the variable $X$ has value $x$ in the unique solution to the equations in $M$ in context $\vec{u}$. Finally, $(M,\vec{u})\models[\vec{Y}\leftarrow\vec{y}]\varphi$ if $(M_{\vec{Y}=\vec{y}},\vec{u})\models\varphi$, where $M_{\vec{Y}\leftarrow\vec{y}}$ is the causal model that is identical to $M$, except that the variables in $\vec{Y}$ are set to $Y=y$ for each $Y\in\vec{Y}$ and its corresponding value $y\in\vec{y}$.
A standard use of causal models is to define actual causation: that is, what it means for some particular event that occurred to cause another particular event. There have been a number of definitions of actual causation given for acyclic models (e.g., (?????????)). In this paper, we focus on what has become known as the modified Halpern–Pearl definition and some related definitions introduced by (?). We briefly review the relevant definitions below. The events that can be causes are arbitrary conjunctions of primitive events (formulas of the form $X=x$); the events that can be caused are primitive events, denoting the output of the model.
Definition 3.1 (Actual cause).
$\vec{X}=\vec{x}$ is an actual cause of $\varphi$ in $(M,\vec{u})$ if the following three conditions hold:
AC1.
$(M,\vec{u})\models(\vec{X}=\vec{x})$ and $(M,\vec{u})\models\varphi$.
AC2.
There is a a setting $\vec{x}^{\prime}$ of the variables in $\vec{X}$, a (possibly empty) set $\vec{W}$ of variables in ${\mathcal{}V}-\vec{X}^{\prime}$, and a setting $\vec{w}$ of the variables in $\vec{W}$ such that $(M,\vec{u})\models\vec{W}=\vec{w}$ and $(M,\vec{u})\models[\vec{X}\leftarrow\vec{x}^{\prime},\vec{W}\leftarrow\vec{w}]\neg{\varphi}$, and moreover
AC3.
$\vec{X}$ is minimal; there is no strict subset $\vec{X}^{\prime}$ of $\vec{X}$ such that $\vec{X}^{\prime}=\vec{x}^{\prime\prime}$ can replace $\vec{X}=\vec{x}^{\prime}$ in AC2, where $\vec{x}^{\prime\prime}$ is the restriction of $\vec{x}^{\prime}$ to the variables in $\vec{X}^{\prime}$.
In the special case that $\vec{W}=\emptyset$, we get the but-for definition. A variable $x$ in an actual cause $\vec{X}$ is called a part of a cause. In what follows, we adopt the convention of Halpern and state that part of a cause is a cause.
The notion of explanation taken from (?) is relative to a set of contexts.
Definition 3.2 (Explanation).
$\vec{X}=\vec{x}$ is an explanation of $\varphi$ relative to a set ${\mathcal{}K}$ of contexts in a causal model $M$ if the following conditions hold:
EX1a.
If $\vec{u}\in{\mathcal{}K}$ and $(M,\vec{u})\models(\vec{X}=\vec{x})\wedge\varphi$, then there exists a conjunct $X=x$ of $\vec{X}=\vec{x}$ and a (possibly empty) conjunction $\vec{Y}=\vec{y}$ such that $X=x\wedge\vec{Y}=\vec{y}$ is an actual cause of $\varphi$ in $(M,\vec{u})$.
EX1b.
$(M,\vec{u}^{\prime})\models[\vec{X}=\vec{x}]\varphi$ for all contexts $\vec{u}^{\prime}\in{\mathcal{}K}$.
EX2.
$\vec{X}$ is minimal; there is no strict subset $\vec{X}^{\prime}$ of $\vec{X}$ such that $\vec{X}^{\prime}=\vec{x}^{\prime}$ satisfies EX1, where $\vec{x}^{\prime}$ is the restriction of $\vec{x}$ to the variables in $\vec{X}^{\prime}$. (This is SC4).
EX3.
$(M,u)\models\vec{X}=\vec{x}\wedge\varphi$ for some $u\in{\mathcal{}K}$.
4 Theoretical Foundations
We approach the first two research questions theoretically.
RQ1
What is a formal rigorous definition of explanation, suitable for image classification?
We view an image classifier (e.g. a neural network) as a probabilistic causal model. Specifically, the endogenous variables are taken to be the set $\vec{V}$ of pixels that the image classifier gets as input, together with an output variable that we call $O$. The variable $V_{i}\in\vec{V}$ describes the color and intensity of pixel $i$; its value is determined by the exogenous variables. The equation for $O$ determines the output of the model as a function of the pixel values. Thus, the causal network has depth 222, with the exogenous variables determining the feature variables, and the feature variables determining the output variable.
We also assume causal independence between the feature variables $\vec{V}$. In our setting, $\vec{V}$ is the set of pixels of the input image, with the values capturing the color and intensity of the pixels. Pixel independence is a common assumption in explainability tools. This is a non-trivial assumption, and it might seem far-fetched, especially if we consider images capturing real objects: if a group of pixels captures, say, a cat’s ear, then a group of pixels below it should capture a cat’s eye. However, we argue that it is, in fact, accurate on images. Indeed, consider the “Photobombing” dataset, described in Section 6.1. This dataset is comprised of partially obscured images, obtained by overlaying random color patches over Imagenet images. An input from this dataset is a perfectly valid image. Indeed, obscuring a part of the input image either by introducing an artificial object or by positioning a real object in front of the primary subject of the classification does not lead to any change in unobscured pixels. For example, obscuring a cat’s ear does not lead to any change in an (unobscured) group of pixels below it, capturing a cat’s eye. Thus, pixel independence holds on general images.
We note that pixels are not concepts: a  [^2] D image is a projection of a scene in the  [^3] D real world on a plane; concepts that are present in the object in a real world can be invisible on this projection, hence the pixel independence.
We refer the reader to (?) for a more in-depth discussion of causal independence between pixel values in image classification. We note that the causal independence assumption is not true in other types of inputs, such as tabular or spectral data. For those types of inputs, assuming independence is clearly an approximation and might lead to inaccurate results. In this paper, however, we focus on images, where causal independence between pixels holds, and it greatly simplifies explanation extraction, as we show below.
Moreover, as the causal network is of depth 222, all parents of the output variable $O$ are contained in $\vec{V}$. Given these assumptions, the probability on contexts directly corresponds to the probability on seeing various images (which the model presumably learns during training).
Given an input image $I$, the set of contexts ${\mathcal{}K}$ that we consider for an explanation is the set ${\mathcal{}K}_{I}$ obtained by all partial occlusions of $I$, where an occlusion sets a part $\vec{Y}$ of the image to a predefined masking color $\vec{y}$. The probability distribution over ${\mathcal{}K}_{I}$ is assumed to be uniform.
Under the assumptions above, the following definition is equivalent to Definition 3.2, as shown by (?).
Definition 4.1 (Explanation for image classification).
An explanation is a minimal subset of pixels of a given input image that is sufficient for the model $\mathcal{N}$ to classify the image, where “sufficient” is defined as containing only this subset of pixels from the original image, with the other pixels set to the masking color.
As our goal is ranking of the pixels according to their importance for the classification, we only consider singleton causes, that is, causes consisting of one pixel and its value. Moreover, for each pixel we only consider two possible values: its original value in the image and a masking color.
Definition 4.2 (Singleton cause for image classification).
For an image $x$ classified by the DNN as $f(x)=o$, a pixel $p_{i}$ of $x$ and its value in $x$ is a cause of $o$ iff there exists a subset $P_{j}$ of pixels of $x$ such that the following conditions hold:
SC1.
$p_{i}\not\in P_{j}$;
SC2.
changing the color of any subset $P^{\prime}_{j}\subseteq P_{j}$ to the masking color does not change the classification;
SC3.
changing the color of $P_{j}$ and the color of $p_{i}$ to the masking color changes the classification.
We call such $P_{j}$ a witness to the fact that $p_{i}$ is a cause of $x$ being classified as $o$.
Definition 4.2 is equivalent to a part of a cause according to Definition 3.1 in binary causal models with all variables being independent of each other. Moreover, under these restrictions, Definition 4.2 is equivalent to the following definition, which is the original definition of an actual cause by (?).
Definition 4.3 (Original definition of actual cause).
$\vec{X}=\vec{x}$ is an actual cause of $\varphi$ in $(M,\vec{u})$ if the following three conditions hold:
AC1.
$(M,\vec{u})\models(\vec{X}=\vec{x})$ and $(M,\vec{u})\models\varphi$.
AC2’.
There exists a partition $(\vec{Z},\vec{W})$ of $\vec{V}$ with $\vec{X}\subseteq\vec{Z}$ and some setting $(\vec{x}^{\prime},\vec{w}^{\prime})$ of the variables in $(\vec{Z},\vec{W})$ such that if $(M,\vec{u})\models Z=z*$ for $Z\in\vec{Z}$, then
  1. 1.
    $(M,\vec{u})\models[\vec{X}\leftarrow\vec{x}^{\prime},\vec{W}\leftarrow\vec{w}^{\prime}]\neg\varphi$. In other words, changing the values of $(\vec{X},\vec{W})$ from $(\vec{x},\vec{w})$ to $(\vec{x}^{\prime},\vec{w}^{\prime})$ changes $\varphi$ from true to false;
  2. 2.
    $(M,\vec{u})\models[\vec{X}\leftarrow\vec{x},\vec{W}\leftarrow\vec{w}^{\prime},\vec{Z}^{\prime}\leftarrow\vec{z}*]\varphi$ for all subsets $\vec{Z}^{\prime}$ of $\vec{Z}$. In other words, setting $\vec{W}$ to $\vec{w}^{\prime}$ should have no effect on $\varphi$ as long as $\vec{X}$ is kept at its current value $\vec{x}$, even if all the variables in an arbitrary subset of $\vec{Z}$ are set to their original values in the context $\vec{u}$.
AC3.
$\vec{X}$ is minimal; there is no strict subset $\vec{X}^{\prime}$ of $\vec{X}$ that satisfies AC1 and AC2’.
The equivalence is based on the following observations. The minimality is satisfied due to only considering singletons. The set $\vec{Z}$ does not play a role due to the independence between the pixels. Removing $\vec{Z}$ and the minimality condition from Definition 4.3 we get Definition 4.2.
Computing causality in image classification is, therefore, NP-complete, based on the above equivalence and the result by (?), who proved NP-completeness of Definition 4.3 for binary causal models. Moreover, following the same reasoning, Definition 4.1 is equivalent to the definition of explanation by (?) for binary causal models, which is DP-complete (?). Hence, there is little hope to find an efficient algorithm for computing exact explanations for image classification.
Finally, the following definition of the degree of responsibility holds for the case of image classifiers and singleton causes as in Definition 4.2.
Definition 4.4 (Simplified responsibility).
[…]
RQ2
Can we compute explanations based on the definition above without opening the black box? What is the complexity of such computation?
[…]
Observation 4.5.
Given an image $x$ and its classification $o$, we can calculate the degree of responsibility of each pixel $p_{i}$ of $x$ by directly applying Definition 4.2, that is, by checking the conditions SC1, SC2, and SC3 for all subsets $P_{j}$ of pixels of $x$ and then choosing a smallest witness subset. While there is an underlying Boolean formula that determines the classification $o$ given the values of the pixels of $x$, we do not need to discover this formula in order to calculate the degree of responsibility of each pixel of $x$.
In the next section we introduce a compositional approach to computing approximate explanations. The main idea is first to rank pixels according to their approximate degree of responsibility, and then use a greedy approach to extract (approximate) explanations. The algorithm for computing an approximate degree of responsibility is based on the notion of a super-pixel $P_{i}$, which is a subset of pixels of a given image. Given an image $x$, we partition it into a small number of superpixels and compute their degree of responsibility for the output of the DNN. Then, we only refine those superpixels with a high responsibility (exceeding a predefined threshold). The scalability of the approach relies on the following observation, which is shown to be heuristically true in our experiments.
Observation 4.6.
The pixels with the highest responsibility for the DNN’s decision are located in super-pixels with the highest responsibility.
Intuitively, the observation holds when pixels with high responsibility do not appear in the superpixels surrounded by other pixels with very low responsibility for the input image classification outcome. While this can happen in principle, we do not encounter this case in practice owing to the continuous nature of images (even when the explanation is non-contiguous). This property is key to the success of our algorithm.
5 Compositional Explanations
Refer to captionCausalRankingAlgorithmRefer to captionExplanationextractionRefer to captionRefer to caption
Figure 2: High level overview of Algorithm 3. Iterating parts 1 and 2 of the algorithm produce a responsibility map (①). The pixels in the image are then ordered by their approximate responsibility and used to produce a minimal sufficient explanation (②), which captures all that the DNN needs to give a classification 'starfish'.
We recall the third research question.
RQ3
Is there an efficiently computable approximation to explanations?
We answer this question affirmatively, by presenting our greedy compositional explanation (CE) algorithm. The general idea is to calculate the responsibility of a superpixel and recursively distribute this responsibility to all pixels within this superpixel. The CE approach in this work consists of the following steps.
  1. 1.
    Given a set of superpixels, compute the responsibility of each superpixel (Section 5.1).
  2. 2.
    Following the responsibility result in Step 1, further refine the superpixel and calculate the responsibility for the refined superpixels (Section 5.2).
We note that the ranking depends on the selected partition. To ameliorate the effect of a particular partition, the algorithm is re-run a number of times with partitions selected independently at random, and the results are averaged across all the partitions (see Section 5.3). Section 5.4 gives a step-by-step example to illustrate the working of the algorithm.
5.1 Computing the responsibility of a superpixel
Given a set of pixels $\mathcal{P}$, we use $\mathds{P}_{i}$ to denote a partition of $\mathcal{P}$, that is, a set $\{P_{i,j}:\bigcup P_{i,j}=\mathcal{P}$ and $\forall j\not=k,P_{i,j}\cap P_{i,k}=\emptyset\}$. The number of elements in $\mathds{P}_{i}$ is a parameter, denoted by $s$; in this work, we consider $s=4$. We refer to $P_{i,j}$ as superpixels. It is insufficient to consider each superpixel in isolation: there is no reason why anyone super pixel should be the cause of a classification. Therefore we create combinations of superpixels from the power set of $\mathds{P}_{i}$, denoted $2^{\mathds{P}_{i}}$.
For an NN $\mathcal{N}$, an input $x$, and a partition $\mathds{P}_{i}$, we can generalize Def. 4.2 to the set of superpixels defined by $\mathds{P}_{i}$. We denote by $r_{i}(P_{i,j},x,\mathcal{N}(x))$ the degree of responsibility of a superpixel $P_{i,j}$ for $\mathcal{N}$’s classification of $x$, given $\mathds{P}_{i}$.
For a partition $\mathds{P}_{i}$, we denote by $X_{i}$ the set of mutant images obtained from $x$ by masking subsets of $2^{\mathds{P}_{i}}$, and by $\tilde{X}_{i}$ the subset of $X_{i}$ that is classified as the original image $x$. Formally,
$\tilde{X}_{i}=\{x_{m}:\mathcal{N}(x_{m})=\mathcal{N}(x)\}.$
We compute $r_{i}(P_{i,j},x,\mathcal{N}(x))$, the responsibility of each superpixel $P_{i,j}$ in the classification of $x$, in Algorithm 1. For a superpixel $P_{i,j}$, we define the set
$\tilde{X}_{i}^{j}=\{x_{m}:P_{i,j}\mbox{ is not masked in }x_{m}\}\cap\tilde{X}_{i}.$
[…]
INPUT: an image $x$, a partition $\mathds{P}_{i}$
OUTPUT: a responsibility map $\mathds{P}_{i}\rightarrow\mathds{Q}$
1:for each $P_{i,j}\in\mathds{P}_{i}$ do
[…]
3:  $r_{i,j}\leftarrow\frac{1}{k+1}$
4:end for
5:return  $r_{i,0},\dots,r_{i,|P_{i}|-1}$
5.2 Iterative refinement of responsibility
Algorithm 1 calculates the responsibility of each superpixel, subject to a given partition. Then, it proceeds with only the high-responsibility superpixels. Note that in general, it is possible that all superpixels in a given partition have the same responsibility. Consider, for example, a situation where the explanation is right in the middle of the image, and our partition divides the image into four quadrants. Each quadrant would be equally important for the classification, hence we would not gain any insight into why the image was classified in that particular way. In this case, the algorithm chooses another partition. It can also be the case that there might be multiple disjoint combinations of high responsibility superpixels in $\mathds{P}_{i}$. Refining them all can be computationally expensive. rex allows the user to choose from various pruning strategies to reduce the amount of work performed.
Different combinations of superpixels may have equal explanatory power. In the event that one superpixel is all that is required, therefore having responsibility 1, the further iterative subdivision of that superpixel is easy. However, if responsibility is split over multiple superpixels, we need a more refined approach. Take, for example, a situation where the right hand side of the image contains the explanation, as can be seen in Figure 6(b). Here responsibility is split over two super pixels. Let us say that the passing combination of superpixels $P_{c}=\{0,2\}$, assuming that we have numbered $P_{i,j}$ consecutively, and that $s=4$. Each superpixel $P_{i,j}$ has responsibility 0.50.50.5. We cannot simply mask superpixel 2 while the refining superpixel 0 as both together are required for the classification. We handle this problem by holding one superpixel to its original value while refining the other and then flipping the procedure. The superpixel held to its original value, but not being refined, does not get additional responsibility. This procedure is relatively scalable to sets containing more than 2 superpixels. In our implementation, rex, $s$ is hard-coded to 444 because this result in powerset of superpixels of manageable size and the refinement process for sets of superpixels just discussed needs to handle at most three superpixels at a time.
[…]
INPUT:   an image $x$ and a partition $\mathds{P}_{i}$
OUTPUT:   a responsibility map $\mathds{P}_{i}\longrightarrow\mathds{Q}$
[…]
2:if $R$ meets termination condition then
3:  return  $R$
4:end if
[…]
6:for each $P_{c}\in(2^{\mathds{P}_{i}}-\emptyset)$ s.t $R(P_{c})\not=0$ do
[…]
8:end for
9:return  $R^{\prime}$
One partition is rarely sufficient for a high-quality explanation, therefore we compute Algorithm 1 many times and compose the results, as shown in Algorithm 2. This calculates the responsibility for each superpixel (Line 1). If the termination condition is met (Lines 2–3), the responsibility map $Q$ is updated accordingly. Otherwise, for each superpixel in $\mathds{P}_{i}$, we refine it and call the algorithm recursively. We use $\cup$ to include these newly computed values in the returned map. The algorithm terminates when: 1) the superpixels in $\mathds{P}_{i}$ are sufficiently refined (containing only very few pixels), or 2) when all superpixels in $\mathds{P}_{i}$ have the same responsibility (this condition is for efficiency).
5.3 Explanation extraction
So far, we assume one particular partition $\mathds{P}_{i}$, which Algorithm 2 recursively refines and calculates the corresponding responsibilities of superpixels in each step by calling Algorithm 1. We note that the choice of the initial partition over the image can affect the values calculated by Algorithm 2, as this partition determines the set of possible mutants in Algorithm 1. We ameliorate the influence of the choice of any particular partition by iterating the algorithm over a set of partitions. Twenty iterations of the algorithm will therefore yield 202020 starting partitions, chosen at random. rex allows the user to choose from a number of different types of random to build these partitions, with uniform being the default.
In Algorithm 3, we consider $N$ initial partitions and compute an average of the degrees of responsibility induced by each of these partitions. In the algorithm, $\mathds{P}^{x}$ stands for a specific partition chosen randomly from the set of partitions, and $r_{p}$ denotes the degree of responsibility of a pixel $p$ w.r.t. $\mathds{P}^{x}$.
[…]
INPUT:   an input image $x$, a parameter $N\in\mathds{N}$
OUTPUT:   an explanation $\mathcal{E}$
[…]
2:for $c$ in 111 to $N$ do
3:  $\mathds{P}^{x}\leftarrow$ sample a partition
[…]
5:  for each $P_{i,j}\in~\mbox{domain of}~R$ do
6:   $\forall p\in P_{i,j}:r_{p}\leftarrow r_{p}+\frac{R(P_{i,j})}{|P_{i,j}|}$
7:  end for
8:end for
[…]
$\mathcal{E}\leftarrow\emptyset$
[…]
12:  $\mathcal{E}\leftarrow\mathcal{E}\cup\{p_{i}\}$
[…]
15:   return  $\mathcal{E}$
16:  end if
17:end for
Algorithm 3 has two parts: ranking all pixels (Lines 1–9) and constructing the explanation (Lines 10–17). The algorithm ranks the pixels of the image according to their responsibility for the model’s output. Each time a partition is randomly selected (Line 3), the iterative responsibility refinement (Algorithm 2) is called to refine it into a set of fine-grained superpixels and calculate their responsibilities (Line 4). A superpixel’s responsibility is evenly distributed to all its pixels, and the pixel-level responsibility is updated accordingly for each sampled partition (Lines 5–7). After $N$ iterations, all pixels are ranked according to their responsibility $r_{p}$.
The remainder of Algorithm 3 follows the method for explaining the result of an image classifier by (?). That is, we construct a subset of pixels $\mathcal{E}$ to explain $\mathcal{N}$’s output on this particular input $x$ greedily. We add pixels to $\mathcal{E}$ as long as $\mathcal{N}$’s output on $\mathcal{E}$ does not match $\mathcal{N}(x)$. This process terminates when $\mathcal{N}$’s output is the same as on the whole image $x$. The set $\mathcal{E}$ is returned as an explanation. We note that this extraction procedure is not normally followed by other familiar XAI tools, as their several different definitions of explanation do not extend to finding minimal sets of pixels which induce the desired classification.
While we approximate the computation of an explanation in order to ensure efficiency of our approach, the algorithm is built on solid theoretical foundations, which distinguishes it from other random or heuristic-based approaches. In practice, while our algorithm uses an iterative average of a greedy approximation, it yields highly accurate results (Section 6). Furthermore, our approach is simple and general, and uses the neural network as a blackbox.
5.4 Illustrative example
Refer to captionRefer to captionRefer to caption① initialpartitionsRefer to captionRefer to captionRefer to caption② refinedpartitionsRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption③ responsibilitymapRefer to caption④ extractedexplanation
Figure 3: The rex algorithm in action: rex creates an initial random partition of an image into 444 sections (①). All combinations of these sections are queried by the model, with further random refinement applied to those sections or combinations of sections that meet the requirements (②). Some combinations, highlighted in green, are classified as bus, others, in red, are not. This results, after several iterations, in a detailed responsibility map (③), from which a minimal passing explanations can be extracted (④).
To illustrate how rex works, consider Figure 3, which is classified as 'bus' by ResNet50, even though there is an occlusion in the middle. Initially, rex picks up an arbitrary partition of the image with four superpixels. This results in 151515 combinations of masking superpixels, though in practice, we only need to examine 141414 combinations as the $15^{th}$ is the entire image, which we have already tested. We use Algorithm 1 to calculate the responsibility of each superpixel. Those superpixels or combinations of superpixels which contribute towards the correct classification of further refined. This iterative refinement is a recursive application of the same initial random partitioning into four superpixels, as in Algorithm 2.
The responsibility mapss shown in Figure 4 demonstrate the importance of Algorithm 3. The initial partitioning of the image can greatly affect the quality of the responsibility map. Figure 4(a) shows the responsibility map after one iteration of the algorithm. While it still indicates well the areas of interest in the image it is also quite rectilinear. Figure 4(b) contains a lot more information about the distribution of responsibility over the pixels in the image. We now have two distinct peaks of responsibility on either side of the central occlusion (a person). Further iterations cause further refinement. By iteration 100 (Figure 4(e)), it is clear that the dominant explanation is around the rear tyre of the bus. The other explanation however, while depressed in relation to the tyre, has not vanished and can still be found through searching the responsibility map (?). Indeed, a disjoint explanation exists for this image, as given in Figure 5(a).
(a) $N=1$
(b) $N=10$
(c) $N=30$
(d) $N=60$
(e) $N=100$
Figure 4: Improvement of rex’s pixel ranking as the number of iterations $N$ increases (Algorithm 3)
(a) Non-contiguous explanation
(b) Contiguous explanation
Figure 5: Two different explanations extracted from the 303030 iteration ranking in Figure 4Figure 5(a) contrasts interestingly with Figure 5(b): it seems that the front tire by itself is not sufficient (a very small rectangular section of the rear tire is included in the explanation), but the rear tire is.
5.5 Termination and complexity
It is clear that the algorithm terminates as soon as the parts can no longer be divided into superpixels. Moreover, the number of calls it performs to the model is linear in the size of the image, as proved in the following lemma.
Lemma 5.1.
The number of calls of  Algorithm 3 to the model is $O(2^{s}nN)$, where $s$ is the size of the partition in each step (in our setting $s=4$), $n$ is the number of pixels in the original image $x$, and $N$ is the number of initial partitions.
Proof.
The computation of responsibilities of superpixels in one partition is $O(2^{s})$, as the algorithm examines the effect of mutating each subset of the superpixels in the current partition. Note that $s$ is a constant independent of the size of the image. The number of steps is determined by the termination condition on the size of a single superpixel, which in the worst case is the same as a single pixel. In our setting, the algorithm terminates when a single superpixel is smaller contains fewer than 101010 pixels111rex also terminates when no combination of active superpixels has the required classification or a user-defined work budget is exceeded.. However, in general, the algorithm can continue down to the level of a single pixel, thus resulting in $n$ pixels in the last step. The algorithm performs $N$ iterations, and every iteration uses a different initial partition. The parameter $N$ is independent of the size of the image. ∎
6 Evaluation
We conduct a large scale investigation into rex, comparing it with a host of popular XAI tools. Our answers to the research questions will largely remain empirical. Just as there is no universally agreed definition of explanation, there is also no single best way to evaluate the quality of an explanation.
RQ4
What are suitable quality measures for explanations? A measure is suitable if it is meaningful with respect to quality and produces results that match our intuition.
A causal explanation is a minimal set of pixels from an image which are sufficient to obtain the same top classification as the complete set of all pixels of an image. This immediately suggests size of explanation as a robust quality measure. As none of the tools (apart from rex) compute minimal, sufficient explanations, we use the mechanism of extraction (Section 5.3) on the output of all tools tested. We also use the same approach to compute insertion and deletion curves over all XAI outputs.
Utility and interpretation of deletion curves
(a) Starfish
(b) Bison
Figure 6: An image with a high deletion curves does not indicate a low quality explanation. The starfish (Figure 6(a)) has a normalized deletion of 0.320.320.32 because their are multiple points of high responsibility in the image. These points indicate the presence of multiple explanations. The bison (Figure 6(b)) has a low deletion of 0.10.10.1 because there is only one point of interest in the image. Deletion is a difficult measure to interpret.
When insertion and deletion curves were introduced in rise (?), the authors stated that a high insertion and a low deletion is preferable. This assertion relies on the assumption that there is only one explanation for the image’s classification, or that the XAI tool finds only one explanation. Moreover, saliency values outside of the single ‘explanation’ must be strictly uninformative: in order to keep deletion low, the saliency map cannot provide information about any substructures present in the image. Pixels must be added back into the image in an order that is essentially random. rex does not work like this: the responsibility map provides detail over the entire pixel space.
As an example, if we examine Figure 6(a), we have an image with a normalized insertion curve of 0.970.970.97 but a normalized deletion curve of 0.320.320.32. There would appear to be tension between the two numbers, as a high insertion suggests a high quality explanation whereas a high deletion suggests a low quality explanation. The mystery is resolved, however, by noting that images can and do contain multiple explanations, as first shown in (?). Indeed, there are four distinct peaks of responsibility in Figure 6(a), each of which corresponds to an independent explanation for starfish (Figure 7).
(a) Explanation 1
(b) Explanation 2
(c) Explanation 3
(d) Explanation 4
Figure 7: The responsibility map rex provides is sufficiently detailed to extract multiple valid explanations for most images. These starfish explanations were extracted from Figure 6(a)).
The deletion score is high because the responsibility map is informative even for the more lowly ranked pixels. In the multiple explanation scenario, one would expect a high deletion score. Uniquely among the tools examined here, rex supports multiple explanations for an image natively (?). In Tables 1, 3 and 5 we follow the usual procedure of indicating the lowest deletion curve value, but stress that lowest does not necessarily indicate best.
We quantify the quality of the explanations with several complementary methods. First, we calculate the minimal explanation size as a percentage of the size of the image. This coincides with Definition 3.2. It also gives an idea of how concentrated the responsibility or saliency maps provided by the XAI tools are. The smaller the area of the explanation, the more the XAI tool was able to isolate precisely those pixels (and only those pixels) required to match the original classification.
Furthermore, we calculate normalized insertion and deletion curves for all images. As stated above, it is not clear how to compare the area under curve (AUC) of a low confidence image against a high confidence image to derive an average AUC performance. It is, after all, not the fault of the XAI tool that the confidence of the model is low on any particular image. To overcome this problem, we follow (?) and normalize all curves by the initial confidence of the model on image under test. Note that this can result in the the insertion curve AUC being greater than 111. This phenomenon occurs when the confidence on the overall image is lower than the confidence of intermediate insertion stages. This makes intuitive sense: if the pixel ranking is accurate, then those pixels that either do not contribute towards the classification, or reduce confidence, are left to be added last. Without these pixels, confidence is potentially much higher.
Comparison with human segmentation
For those datasets that contain segmentation information, we also measure the intersection of the minimal explanation with the segmentation. This segmentation is human-provided and we argue corresponds most closely with what a human considers acceptable. On VOC2012, for example, we would like an minimal explanation to reside mostly inside the human-provided segmentation mask and therefore contain only a small number of pixels outside the mask. For datasets that feature occluded objects, a good explanation would have as few pixels from inside the occlusion as possible, as the occlusion should not make a contribution towards the model’s classification.
To answer RQ4, we argue that the size of explanations, as well as insertion curves and the intersection with the occlusion are measures that match our intuition. However, we argue that deletion curves are not, as they remain high in presence of multiple explanations.
6.1 Experimental Setup
We answer research questions RQ5–RQ7 experimentally. We have implemented the proposed explanation approach in the publicly available tool rex. For all other XAI tools, apart from rise, we used the Captum PyTorch library (?) [^2] https://www.captum.ai. For rise, we used the authors’ implementation [^3] https://github.com/eclique/RISE, which we lightly altered to use PyTorch rather than tensorflow.
In the evaluation, we compare rex with a wide range of explanation tools, specifically Grad-CAM (?), KernelShap (?), GradientShap (?), rise (?), lime (?), ig (?), noisetunnel (?). For the ResNet50 model only, we also include lrp. We were unable to make the captum version of lrp run on either the ViT or the ConvNext model.
While our main interest is in black-box explainability, we have included methods which require access to the model gradient, such as gradientshap and ig, and internal model layers, i.e. Grad-CAM. We call those tools which require internal access to the model ‘white-box’; those that require the access to the gradient we term ‘grey-box’, and those that just need the model’s classification, we call ‘black-box’. rex is a black-box method. PyTorch permits optional access to the gradient, so tools such as ig require at least some flexibility when calling the model. This is not guaranteed with proprietary models: hence our grey-box category.
We use 444 data sets: ImageNet-1k-mini validation (?), ECSSD (?), Pascal VOC2012 (?), and a “Photobombing” dataset we created. Imagenet1k-mini comes with labels for ground truth. VOC2012 has labels and segmentation data. ECSSD comprises complex images which come with a human-provided segmentation. We created the “Photobombing” dataset by inserting black occlusions into ImageNet images, meaning we have the original label and also know the pixel coordinates of the occlusions. These occlusions are placed so as not to change the model classification. We use the torchvision [^4] https://pytorch.org/vision/stable/index.html implementations of ResNet50, ViT and ConvNext-Large with default weights on all data sets.
All experiments were conducted using a server running Ubuntu 20.04 with an Nvidia A40 GPU. Moreover, all tools were used with default settings. For rex in particular, this means that it performs 20 iterations of the algorithm ($N$ in Algorithm 3), with a minimum superpixel size of 10 pixels. The pruning strategy for the work queue is “area” (the passing mutants are ordered by size) and only one item is kept in the work queue at a time, so we only refine the smallest passing mutant. Partitions are created uniformly at random. Other strategies and partitioning distributions are available in the configuration.
rex uses four superpixels per partition. This is practical for images as we can split the image with just one call to a random number generator. The number of mutants produced is also relatively small (151515), which can usually be fit into a single batch for model processing. Having more initial superpixels does not lead to greater expressivity: we iteratively refine passing combinations of superpixels further, essentially recreating a more detailed initial partition produced by a greater number of superpixels. The termination conditions for the partition refinement in Algorithm 2 are: 1) the area of a superpixel is less than 101010 pixels of (resized) input image, 2) the four superpixels share the same responsibility or 3) no combination of superpixels produces the desired classification.
6.2 Results
RQ5
What is the precision of explanations computed by our algorithm compared to other XAI methods?
RQ6
Is there a trade-off between precision and compute cost of the explanations?
RQ7
Can black-box methods achieve the same quality of explanations as white and grey-box methods?
Table 1: Results for ResNet50 model.
(a) Results on Voc
tool area ($\downarrow$) std ins ($\uparrow$) del IN ($\uparrow$) OUT ($\uparrow$)
Grad-CAM 0.0832 0.1202 1.0488 0.2756 0.3202
noisetunnel 0.2867 0.2627 0.6953 0.0851 0.3241
lrp 0.3936 0.2828 0.5267 0.2953 0.2093
ig 0.4904 0.2768 0.4899 0.1228 0.2215
GradientShap 0.3894 0.3002 0.5935 0.0651 0.2872
KernelShap 0.7741 0.2084 0.188 0.1941 0.1636
lime 0.1124 0.1616 0.963 0.2329 0.3039
rex 0.0427 0.0505 1.2218 0.2148 0.6432
rise 0.1271 0.1547 0.9358 0.2264 0.2861
(b) Results on Imagenet
tool area ($\downarrow$) std ins ($\uparrow$) del IN ($\uparrow$) OUT ($\uparrow$)
Grad-CAM 0.0486 0.0839 1.0059 0.4073
noisetunnel 0.1847 0.2131 0.7612 0.1274
lrp 0.2749 0.2418 0.5972 0.3567
ig 0.4063 0.264 0.534 0.1767
GradientShap 0.2861 0.2696 0.6474 0.0983
KernelShap 0.6686 0.2472 0.2549 0.2649
lime 0.0649 0.0969 0.9871 0.3313
rex 0.0333 0.0351 1.1101 0.3403
rise 0.1026 0.1254 0.8898 0.3728
(c) Results on “Photobombing”
tool area ($\downarrow$) std ins ($\uparrow$) del IN ($\uparrow$) OUT ($\uparrow$)
Grad-CAM 0.0455 0.0642 0.9843 0.3625 0.8505
noisetunnel 0.1565 0.1832 0.7604 0.1348 0.8732
lrp 0.2429 0.205 0.6049 0.3687 0.9058
ig 0.4184 0.2311 0.5178 0.1853 0.8806
GradientShap 0.296 0.2373 0.637 0.114 0.8533
KernelShap 0.6601 0.2246 0.2527 0.2612 0.9188
lime 0.0507 0.0523 0.9937 0.3046 0.8921
rex 0.0297 0.026 1.0481 0.2947 0.9739
rise 0.0787 0.0942 0.8906 0.2977 0.8798
[…]
(a) Results on ECSSD
tool area ($\downarrow$) std ins ($\uparrow$) del IN ($\uparrow$) OUT ($\uparrow$)
Grad-CAM 0.0836 0.1288 1.0226 0.2915 0.6654
noisetunnel 0.2695 0.276 0.7194 0.119 0.5955
lrp 0.3545 0.292 0.5721 0.305 0.3418
ig 0.4437 0.3016 0.5552 0.1377 0.4187
GradientShap 0.3525 0.3182 0.6444 0.0977 0.5379
KernelShap 0.7257 0.2403 0.2348 0.2369 0.244
lime 0.1013 0.1579 1.0024 0.2404 0.6538
rex 0.0423 0.0522 1.2176 0.2214 0.6205
rise 0.1165 0.1599 0.95 0.2752 0.5451
Tables 1, 3 and 5 show the experimental results of evaluation of rex against 888 other XAI tools over 444 datasets, with three different models. The columns in the tables are: mean area (area), $\sigma$ of area (standard deviation), normalized insertion curves AUC (ins), normalized deletion curves AUC (del), proportion of explanation inside the relevant segment (IN), and the fraction of the explanation that is outside the relevant segment (OUT), shown only where appropriate. Bold indicates best result in category for all columns. Note that while we have indicated lowest AUC for deletion, lower here is not necessarily better, as we argued earlier.
Resnet
Table 1 shows the results with the ResNet50 model. rex consistently produces the most precise explanations. This is reinforced by the observations that rex also produces the highest insertion AUC of any of the tools. rex does particularly well on the ‘Photobombing‘ dataset, where its explanations are almost entirely (0.970.970.97) disjoint with the inserted occlusion (Figure 12). Figure 8 shows that rex has the most concentrated explanations and consistent results, with very few outliers. The grey-box tools perform worse than the other methods. Note that rex even outperforms Grad-CAM.
Table 3: Results for ViT model.
(a) Results on Voc
tool area ($\downarrow$) std ins ($\uparrow$) del IN ($\uparrow$) OUT ($\uparrow$)
Grad-CAM 0.7148 0.181 0.2466 0.1889 0.1727
noisetunnel 0.4988 0.3019 0.4565 0.0633 0.2315
ig 0.5652 0.2632 0.4274 0.0943 0.1946
GradientShap 0.5021 0.2992 0.4761 0.0551 0.2267
KernelShap 0.669 0.2603 0.2579 0.2579 0.1686
lime 0.277 0.2218 0.7191 0.112 0.26
rex 0.2206 0.1711 0.6985 0.1377 0.5496
rise 0.4049 0.2491 0.5986 0.097 0.2365
(b) Results on Imagenet
tool area ($\downarrow$) std ins ($\uparrow$) del IN ($\uparrow$) OUT ($\uparrow$)
Grad-CAM 0.6277 0.203 0.2943 0.2538
noisetunnel 0.3903 0.2711 0.5341 0.098
ig 0.4706 0.2505 0.4833 0.1481
GradientShap 0.3871 0.2691 0.5586 0.0916
KernelShap 0.5483 0.2651 0.3541 0.3519
lime 0.2246 0.1937 0.7264 0.1668
rex 0.1689 0.133 0.687 0.2015
rise 0.3356 0.2333 0.5997 0.1664
[…]
(a) Results on “Photobombing”
tool area ($\downarrow$) std ins ($\uparrow$) del IN ($\uparrow$) OUT ($\uparrow$)
Grad-CAM 0.6025 0.1907 0.303 0.2715 0.9196
noisetunnel 0.3647 0.2447 0.5439 0.1012 0.8835
ig 0.4598 0.2251 0.4806 0.1486 0.8889
GradientShap 0.3726 0.2465 0.5607 0.0967 0.8802
KernelShap 0.5168 0.233 0.37 0.3665 0.9186
lime 0.1748 0.1458 0.7365 0.1778 0.8891
rex 0.1402 0.1046 0.6883 0.2073 0.9551
rise 0.2803 0.2056 0.6119 0.1741 0.8915
(b) Results on ECSSD
tool area ($\downarrow$) std ins ($\uparrow$) del IN ($\uparrow$) OUT ($\uparrow$)
Grad-CAM 0.6741 0.2181 0.2731 0.2344 0.2494
noisetunnel 0.4298 0.3001 0.5371 0.0731 0.4296
ig 0.4727 0.2723 0.5184 0.1082 0.3409
GradientShap 0.4179 0.2929 0.5772 0.0685 0.4134
KernelShap 0.6192 0.2729 0.321 0.3188 0.2466
lime 0.2698 0.2319 0.7392 0.123 0.5548
rex 0.2008 0.1653 0.7286 0.1475 0.5334
rise 0.3687 0.2585 0.6597 0.1178 0.4468
ViT
Table 3 shows the results with the ‘vit_b_32‘ model from torchvision. Explanations across all tools are much larger than for ResNet50model, which seems to be a feature of the model. rex, however, still manages to produce explanations with the lowest number of redundant pixels. In general, rex, lime and Grad-CAM are the best performing tools. Even in cases where lime slightly outperforms rex (e.g., on ECSSD IN), with 0.730.730.73 of its explanation inside the segmentation against 0.720.720.72, the effect is very small. Figure 9 shows that rex again has the most concentrated explanations and consistent results, with very few outliers. lime is the closest comparable tool, though it has many more outliers and a higher median value. The significant difference in explanation size between the ResNet model and ViT model is, perhaps, of independent interest.
Table 5: Results for convnext model.
(a) Results on Voc
tool area ($\downarrow$) std ins ($\uparrow$) del IN ($\uparrow$) OUT ($\uparrow$)
Grad-CAM 0.0556 0.0948 0.8635 0.2497 0.3345
noisetunnel 0.1972 0.2443 0.689 0.117 0.3614
ig 0.4152 0.2565 0.465 0.2387 0.2025
GradientShap 0.2268 0.2609 0.6411 0.0998 0.3327
KernelShap 0.5739 0.3001 0.2926 0.2924 0.1678
lime 0.0689 0.0972 0.8096 0.2908 0.3042
rex 0.0361 0.038 0.876 0.2057 0.6295
rise 0.0759 0.1129 0.8136 0.3005 0.2863
(b) Results on Imagenet
tool area ($\downarrow$) std ins ($\uparrow$) del IN ($\uparrow$) OUT ($\uparrow$)
Grad-CAM 0.0337 0.0553 0.8613 0.4413
noisetunnel 0.1116 0.1608 0.7682 0.1886
ig 0.3414 0.2168 0.4993 0.3471
GradientShap 0.1394 0.183 0.706 0.1678
KernelShap 0.4357 0.2855 0.4042 0.4061
lime 0.0493 0.0707 0.8228 0.4747
rex 0.0287 0.0236 0.8788 0.3487
rise 0.0615 0.084 0.8058 0.5136
(c) Results on “Photobombing”
tool area ($\downarrow$) std ins ($\uparrow$) del IN ($\uparrow$) OUT ($\uparrow$)
Grad-CAM 0.0426 0.0527 0.8584 0.3688 0.8695
noisetunnel 0.0768 0.0925 0.7916 0.2077 0.8755
ig 0.3144 0.1731 0.5454 0.3847 0.8783
GradientShap 0.1335 0.1351 0.7095 0.2186 0.85
KernelShap 0.3829 0.2405 0.4683 0.468 0.9183
lime 0.0513 0.0551 0.8643 0.4161 0.8947
rex 0.0296 0.0236 0.8475 0.3173 0.9748
rise 0.063 0.0723 0.8129 0.4769 0.9015
[…]
(a) Results on ECSSD
tool area ($\downarrow$) std ins ($\uparrow$) del IN ($\uparrow$) OUT ($\uparrow$)
Grad-CAM 0.0574 0.1023 0.8865 0.2694 0.6846
noisetunnel 0.2164 0.2734 0.6424 0.1067 0.5913
ig 0.421 0.2925 0.4202 0.2102 0.363
GradientShap 0.2434 0.284 0.5911 0.0987 0.5671
KernelShap 0.5616 0.3152 0.2898 0.2977 0.2468
lime 0.0849 0.1377 0.7157 0.3183 0.5841
rex 0.0364 0.0405 0.9012 0.2077 0.5722
rise 0.0823 0.1387 0.7996 0.3073 0.5045
ConvNext
Table 5 shows the results with the ‘convnext_large‘ model from torchvision. rex, lime, Grad-CAM and rise are the best performing tools. It is interesting to note that 333 of these tools fall into our ‘black-box‘ category. Figure 10 shows that rex has the best performance again, this time slightly ahead of Grad-CAM, as Grad-CAM has many more outliers than rex and a slightly larger inter-quartile range. Explanation size is comparable with that of the ResNet model, making the ViT model somewhat of an outlier.
Figure 8: Box plot of all tools over all datasets with a ResNet50 model. rex has the lowest median value and also the smallest number of outliers.
Figure 9: Box plot of all tools over all datasets with a ViT model. rex has the lowest median value and also the smallest number of outliers. ViT explanations are noticeably larger than those for ResNet50 and ConvNext.
Figure 10: Box plot of all tools over all datasets with a ConvNext model. rex has the lowest median value and also the smallest number of outliers. It even outperforms Grad-CAM, a purely white-box tool.
Runtime
We measured the running time of all tools. As all tools were evaluated on the same hardware with the same models and datasets, the results should be consistent. The white and gray-box methods are clearly the fastest, as expected. Indeed, Grad-CAM is essentially instantaneous, as the overhead over one call to the model is insignificant. No black-box tool can compete with this, by the nature of black-box tools. All the gradient based methods took 3$\approx$ 3≈ 3 seconds to produce initial maps. While they are still fast, the quality of their explanations is lower.
Both rex and lime took 4$\approx$ 4≈ 4s to produce maps, with rise being much slower (10$\approx$ 10≈ 10s). If we take into account the extra time for the explanation extraction (which only rex does by default), the efficiency picture changes slightly. rex’s total compute time increases only slightly (to 6$\approx$ 6≈ 6s) due the the uniformly small size of its explanations. Other tools suffer more though, even Grad-CAM’s average compute time increases to 4$\approx$ 4≈ 4s. rex is an efficient black-box tool: for RQ6 therefore, our results show that there is little trade-off between quality and compute cost.
6.3 Discussion
Broadly speaking, the tools can be split into three different sets. The gradient-based methods perform least well on all metrics, including overlap with a human-provided segmentation mask. The explanations they provide are diffuse and large. noisetunnel uses ig as its primary saliency method, and in return for a relatively small computational overhead, noisetunnel greatly improves the explanation quality over the ig baseline.
Figure 11 shows cumulative plots of explanation sizes over the different tested models. Again, rex dominates the other curves. lime and Grad-CAM are, in general, the next best performing tools. The poor performance (almost linear) of KernelShap is interesting, suggesting that image explainability is not a natural fit for this method.
(a) ResNet50
(b) ViT
(c) ConvNext
Figure 11: Cumulative plots of explanations sizes with 333 different models over all datasets. rex consistently produces smaller explanations, meaning it identifies precisely those pixels required for the classification, with fewer unnecessary pixels.
lime, when used for image classification, works best when provided with a segmentation mask of the image. For common datasets such as the ones we investigate, many high quality algorithms exist. rex does not require such existing segmentation masks. (?) conducted an investigation on brain MRI explainability using a different model and different data set and a similar, though not overlapping, set of explainability tools. On this dataset, lime’s performance is hampered by medically meaningless segmentation. They found that rex did not suffer from this problem, but performed well compared to the other tools on this different model and data.
The performance of rex is notable across all datasets for being very consistent in terms of the evaluation metrics. The other tools examined are not as consistent. Note that the values for IN and OUT in Tables 1, 3 and 5 are percentages of the explanation contained (or outside) the segmentation. As rex’s explanations are already smaller than the others, this means that, in absolute terms, rex also has fewer extraneous pixels. This holds because other tools (such as lime) may have similar IN values, but a larger explanation size.
Figure 12: Photo Bombing images and output from rex
We evaluated the results produced by rex with those produced by other explainability tools on a number of standard measures proposed in the literature, such as the size of explanations, insertion and deletion curves. For the datasets containing the ground truth or some part of it, such as the segmentation information or the partial occlusion, we also measured the percentage of the explanation that is clearly outside of an area where it should be contained (such as the occlusion in partially occluded areas). For all other measures, rex shows superior performance, which answers RQ5.
Finally, for RQ7, we did not exhaustively compare the results against all possible explainability methods. Rather, we chose the state-of-the-art primary attribution methods from the Captum library.
7 Conclusions and Future Work
In this paper we described an approach to explainability of image classifiers that is rooted in actual causality and computes explanations based on the formal definitions of sufficient explanations. As the exact computation is intractable, we described a modular algorithm for computing approximate explanations using a causal ranking of parts of the input image, viewing the classifier as a black box. Our experiments demonstrate that an implementation of our algorithm in the tool rex produces results that are superior to those of state-of-the-art tools according to standard measures.
In the future, we will apply rex to other modalities that lend themselves to occlusion-based reasoning, such as tabular data, time series, spectroscopy, etc. We will also explore different domains of application, where precise explanations are instrumental, in particular mission-critical and safety-critical domains, such as the healthcare AI and autonomous vehicles. The healthcare domain will require some adaptations of our algorithms, since the quality of medical images is different from the quality of general images (e.g., there are fewer colors in an X-ray than in a general image). In the automotive domain, we will adapt rex to analyze object detectors and also devise new algorithms that work in (near) real-time, to address the needs of explainability for autonomous vehicles.

### v3

@@
 Causal Explanations for Image Classifiers
-Abstract
-Existing algorithms for explaining the output of image classifiers use different definitions of explanations and a variety of techniques to extract them.
-However, none of the existing tools use a principled approach based on formal definitions of causes and explanations for the explanation extraction.
+Abstract.
+Existing algorithms for explaining the output of image classifiers use different definitions of explanations and a variety of techniques to find them.
+However, none of the existing tools use a principled approach based on formal definitions of cause and explanation.
 In this paper we present a novel black-box approach to computing explanations grounded in the theory of actual causality.
@@
 We prove termination of our algorithm and discuss its complexity and the amount of approximation compared to the precise definition.
-We implemented the framework in a tool rex and we present experimental results and a comparison with state-of-the-art tools.
-We demonstrate that rex is the most efficient tool and produces the smallest explanations, in addition to outperforming other black-box tools on standard quality measures.
-1 Introduction
+We implemented the framework in a tool, rex, and we present experimental results and a comparison with state-of-the-art tools.
+We demonstrate that rex is the most efficient black-box tool and produces the smallest explanations, in addition to outperforming other black-box tools on standard quality measures.
+copyright: ccjournalvolume: 0article: 0publicationmonth: 0journalyear: 0
+1. Introduction
 (a) Ladybug
@@
 (d) Explanation
-[…]
+Figure 1.
+A ladybug (1(a)), its responsibility map (1(b)), the heat map (1(c)), which is a projection of the responsibility map on a plane overlaid on the original image, and a causal explanation (1(d)).
+The minimal causal explanation computed by our tool rex is less than $1\%$ of the image.
 Neural networks (NNs) are now a primary building block of many computer vision systems.
-NNs are complex non-linear functions with algorithmically generated (and not engineered) coefficients.
-In contrast to traditionally engineered image processing pipelines it is difficult to retrace how the pixel data are interpreted by the layers of the network.
+NNs are complex non-linear functions with algorithmically generated coefficients.
+In contrast to traditionally engineered image processing pipelines, it is difficult to retrace how the pixel data are interpreted by the layers of the network.
 Moreover, in many application areas, in particular healthcare, the networks are proprietary, making the analysis of the internal layers impossible.
-This “black box” nature of neural networks creates demand for techniques that explain why a particular input yields the output that is observed without understanding the model’s parameters and their influence on the output.
+The “black box” nature of neural networks creates demand for eXplainable AI (XAI) techniques that show why a particular input yields the observed output without understanding the model’s parameters and their influence on that output.
 An explanation of an output of an automated procedure is essential in many areas, including verification, planning, diagnosis and the like.
-A good explanation can increase a user’s confidence in the result.
+An informative explanation may alter a user’s confidence in the result.
 Explanations are also useful for determining whether there is a fault in the automated procedure: if the explanation does not make sense, it may indicate that the procedure is faulty.
-It is less clear how to define what a good explanation is.
-There have been a number of definitions of explanations over the years in various domains of computer science (???), philosophy (?) and statistics (?).
+As argued by bhusalface, it is important to recognize that a “good” explanation should reveal what information a model is using to make its decision, even if that information is not intuitive to a human.
+There have been a number of definitions of explanations over the years in various domains of computer science (CH97; Gar88; Pea88), philosophy (Hem65) and statistics (Sal89).
 Black-box explanations for the results of image classifiers are typically based on or are given in the form of a ranking of the pixels, which is a numerical measure of importance: the higher the score, the more important the pixel is for the NN’s classification outcome.
 Often these rankings are presented in a form of a heat map, with higher scores corresponding to hotter areas.
-A user-friendly explanation can then be a subset of highest-ranked pixels that is sufficient for the original classification outcome.
 This paper addresses the following research questions:
-RQ1
-What is a formal rigorous definition of explanation, suitable for image classification?
-RQ2
+RQ1:
+What is a formal, rigorous definition of explanation, suitable for analyzing image classifiers?
+RQ2:
 Can we compute explanations based on the definition above without opening the black box? What is the complexity of such computation?
-RQ3
-Is there an efficiently computable approximation to explanations?
-RQ4
-What are suitable quality measures for explanations? A measure is suitable if it is meaningful with respect to quality and produces results that match our intuition.
-RQ5
-What is the quality of explanations computed by our algorithms compared to other black-box methods?
-RQ6
+RQ3:
+Is there an efficiently computable approximation for explanations?
+RQ4:
+What are suitable quality measures for explanations?
+RQ5:
+What is the quality of explanations computed by our algorithms compared to other XAI methods?
+RQ6:
 Is there a trade-off between quality and compute cost of the explanations?
-RQ7
+RQ7:
 Can black-box methods achieve the same quality of explanations as white-box methods?
-[…]
-We compare the experimental results of our implementation of the algorithm in a tool rex with those of other black-box tools on a number of benchmark datasets: ImageNet, VOC2012, and a set of partially occluded images.
-Our results show that rex is on par with the most efficient tools in terms of execution time, while producing the smallest explanations.
-We also analyse the explanations using the standard measures of insertion and deletion curves and the intersection with the background or the occlusion.
-The results show that rex outperforms the other tools on insertion curves and the overlap with irrelevant parts of the image.
-We discuss the logic behind deletion curves and demonstrate that low deletion is a measure of the quality of the model for particular types of images, rather than for the quality of an explanation.
-We also argue that explanations computed by rex are so small that any further reduction in their size is meaningless.
+Our algorithms are based on the formal definition of explanation in the theory of actual causality by Hal19 and its adaptation to image classifiers by CH24.
+Essentially, an explanation in the context of image classifiers is a minimal subset of pixels with their original values that is sufficient for obtaining the same classification as the original image, given that the rest of the image has its values replaced with data from a predefined dataset.
+We view the task of image classification as a causal model (Section 3) of depth two, with the classification itself computed in one internal node, thus capturing the opacity of the network.
+The algorithm (Section 5) then calculates an approximate explanation, where the approximation is in the size of the explanation (Figure 1).
+That is, an explanation computed by our algorithm is sufficient for the classification, but it might not be minimal, though it is very close to minimal on real images: in our experiments, explanations average between $3\%-5\%$ of the image, which is smaller than any of the other black-box tools.
+Size is a useful metric here as it indicates the degree of extraneous information contained in the explanation.
+We formally prove termination and complexity of our algorithm and discuss its approximation ratio.
+We compare the experimental results of our implementation of the algorithm in a tool, rex, with those of other XAI tools, both white-box and black-box, on a number of benchmark datasets: ImageNet, VOC2012, ECSSD, and a set of partially occluded images.
+Our results show that rex is on par with the most efficient black-box tools in terms of execution time, while producing the smallest explanations.
+We also analyze the explanations using the standard measures of insertion and deletion curves (petsiuk2018rise) and the intersection with the background or the occlusion.
+The results show that rex outperforms the other tools on insertion curves and on the overlap with irrelevant parts of the image.
+We discuss the logic behind deletion curves and demonstrate that low deletion is a measure of the quality of the model for particular types of images, rather than for the quality of an explanation (Section 6.3).
+We also argue that explanations computed by rex are so small that any further reduction in their size is not useful (kelly2025big).
 The rest of the paper is organized as follows.
@@
 The tool rex is open source and available at https://github.com/ReX-XAI/ReX.
-All benchmark sets are publicly available and referenced in Section 6.
-2 Related Work
-There is a large body of work on explaining the results of image classifiers.
-As our primary interest is in black-box approaches, we only give a brief overview of the white-box algorithms here.
-The existing approaches to post hoc explainability can be largely grouped into two categories: propagation and perturbation.
-Propagation-based explanation methods are often regarded as more efficient.
+All models and benchmark sets are publicly available and referenced in Section 6.
+2. Related Work
+The taxonomy of XAI methods and styles is large (schwalbe2024comprehensive).
+Our method and tool, rex, produces local, post hoc, black-box causal explanations of images.
+By local, we mean that the explanation applies only to one image, not to all images with the same classification.
+This is in contrast to a global explanation, which explains an entire class.
+Post hoc means that the method is applied to a model which is not inherently explainable and which has not been designed with explainability in mind.
+By black-box, we mean that there is no access at all to model architecture, gradients, or internals.
+An XAI method with free access to any layer of the model is white-box.
+An XAI tool with access to discretionary information, such as the gradient, we term grey-box.
+A proprietary system need not expose any information to the user other than the top-$1$ class.
+As our primary interest is in black-box, post hoc approaches, we only give a brief overview of other algorithms here.
+The existing approaches to post hoc explainability can largely be grouped into two categories: propagation and perturbation.
+Propagation-based explanation methods are usually more computationally efficient.
 They back-propagate a model’s decision to the input layer to determine the weight of each input feature for the decision.
-Grad-CAM by (?) only needs one backward pass and propagates the class-specific gradient into the final convolutional layer of a DNN to coarsely highlight important regions of an input image.
-Guided Back-Propagation (gbp) by (?) computes the single and average partial derivatives of the output to attribute the prediction of a DNN.
-Integrated Gradients (ig) by (?) further uses two axioms called sensitivity and completeness for the problem of how to attribute the classification by a deep network to its input features.
-noisetunnel by (?) adds gaussian noise to the image before applying another XAI method.
-The noise serves to smooth any resultant saliency map, improving explanation quality.
-shap (SHapley Additive exPlanations) by (?) goes beyond propagation and identifies inputs that are similar to the input for which the output is to be explained, and ranks the features of the input according to their difference.
-A key advantage of shap is that it does not require counterfactuals.
-gradientshap is gradient-based method that approximates Shapley values via gradient integration and a noise tunnel.
-Both ig and gradientshap require access to model gradients while computing explanations.
-They form part of a family of whitebox explanation methods which differ from the Grad-CAM family in that they do not rely on examination of specified activation layers in the model.
-Other methods for approximating Shapley values are numerous, among which KernelShap (?) and DeepLiftShap (?) are probably the most used.
-KernelShap uses the lime framework to approximate Shapley values, whereas DeepListShap uses the DeepLift (?) method to the same end.
-In contrast to propagation-based explanation methods, perturbation-based explanation approaches explore the input space directly in search for an explanation.
-The exploration/search often requires a large number of inference passes, which incurs significant computational cost when compared to propagation methods.
-There are many sampling methods, most of which are based on random search or heuristics.
-The chief advantage of perturbation-based methods is that they are model independent and do not require any access to model internals.
-Given a particular input, lime by (?) samples the the neighborhood of this input and creates a inherently explainable model to approximate the system’s local behavior; owing to the high computational cost of this approach, the ranking uses super-pixels via segmentation instead of individual pixels.
-(?) replace the natural distribution of the input by a user-defined distribution and then use the Shapley Value method to analyze combinations of input features and to rank their importance.
-(?) estimate the importance of input features by measuring the the flow of information between inputs and outputs.
-Both the Shapley Value and the information-theoretic approaches are computationally expensive.
-In rise by (?), the importance of a pixel is computed as the expectation over all local perturbations conditioned on the event that the pixel is observed.
-(?) have introduced the concept of “extreme perturbations” to improve the perturbation analysis in the extremal algorithm.
-More recently, spectrum-based fault localisation (SBFL) has been applied to explaining image classifiers.
-The technique has been implemented in the tool deepcover by (?).
-The method and tool, rex, presented in this paper is a perturbation-based approach and addresses the limitations of existing black-box methods in two aspects.
-The feature masking in rex uses causal reasoning that provides a guarantee (subject to the assumption).
-Owing to its guided iterative refinement, rex is computationally efficient, while still producing explanations that are the state of the art.
-3 Background on Actual Causality
-In this section, we briefly review the definitions of causality and causal models introduced by (?) and relevant definitions of causes and explanations in image classification by (?).
-The reader is referred to (?) for further reading.
+Grad-CAM (CAM) only needs one backward pass and propagates the class-specific gradient into the final convolutional layer of a DNN to coarsely highlight important regions of an input image.
+Guided Back-Propagation (gbp(springenberg2015striving) computes the single and average partial derivatives of the output to attribute the prediction of a DNN.
+lrp (LRP) uses layer-wise back-propagation applied to all layers of the model.
+The model’s output is decomposed into different degrees of relevance by applying different rules for each layer.
+In contrast to propagation-based explanation methods, perturbation-based explanation approaches explore the input space directly in search of an explanation.
+The exploration/search often requires a large number of inference passes, which may incur significant computational cost, especially when compared to propagation methods.
+There are many ways to create these perturbations, most of which are based on random search or heuristics.
+The chief advantage of perturbation-based methods is that they are model independent and do not, in general, require any access to model internals.
+Integrated Gradients (ig(sundararajan2017axiomatic) relies only on the model’s classification and gradient.
+As a model does not need to expose the gradient, we consider ig a grey-box method.
+It progressively introduces pixels from an image onto an underlying baseline value and computes an explanation based on the effect of these introductions on the output gradient.
+shap (SHapley Additive exPlanations) (lundberg2017unified) is a family of different techniques, all of which seek to estimate Shapley values (winter2002shapley).
+Shapley values measure the contribution, either positive or negative, of a feature towards an outcome.
+gradientshap is a gradient-based method that approximates Shapley values via gradient integration, in a similar fashion to ig.
+There are numerous other methods for approximating Shapley values, among which KernelShap (lundberg2017unified) and DeepLiftShap (lundberg2017unified) are probably the most used.
+DeepLiftShap uses the DeepLift (deeplift) method whereas KernelShap uses the lime framework to approximate Shapley values,
+Given a particular input, lime (lime) samples the neighborhood of a given input and builds an inherently explainable model to approximate the system’s local behavior.
+An inherently explainable model for lime is usually a variant on some linear regression model.
+Owing to the high computational cost of this approach, lime relies on a separate segmentation algorithm to break an image into superpixels, that is, clusters of connected pixels that share similar characteristics, such as color, intensity, or texture.
+The quality of this initial segmentation is vital for good performance, as lime does not refine it further (Knab2024BeyondPE).
+A lime explanation is therefore necessarily tied to the provided segmentation.
+When this segmentation does not align with relevant features, lime’s performance suffers (blake2024explainable).
+For instance, if the segmentation is too large or too crude, lime’s explanations contain irrelevant or unnecessary pixels.
+In rise (petsiuk2018rise), the importance of a pixel is computed as the expectation over all local perturbations conditioned on the event that the pixel is observed.
+Rather than relying on an initial segmentation, rise creates random occlusions of the input image.
+More recently, spectrum-based fault localization (SBFL) has been applied to explaining image classifiers.
+The technique has been implemented in the tool deepcover (sun2020explaining).
+While not an independent XAI method, noisetunnel (noisetunnel) adds gaussian noise to the image before applying another XAI method, usually ig.
+The noise serves to smooth the resultant saliency map.
+None of the methods mentioned above agree on what constitutes an explanation.
+lime produces a locally explainable model, but the user is usually presented with a heat map of the local model’s weights.
+rise also produces a heat map, but its output is not directly comparable with lime.
+gradientshap output is also usually visualized as a heatmap, but with both positive and negative values.
+rex, the method and tool presented in this paper, is a perturbation-based approach and addresses the limitations of existing black-box methods in two aspects.
+rex does not rely on an initial segmentation as does lime, nor does it simply rely on unguided random occlusions, as does rise.
+The feature masking in rex uses causal reasoning to guide the creation of occlusions.
+rex output is also different from other methods: rex shows the user which pixels are sufficient for a classification.
+rex tests this sufficiency against the model itself, unlike lime for example, which uses a locally trained model.
+Owing to its guided iterative refinement, rex is computationally efficient, while still producing explanations that are state of the art.
+3. Background on Actual Causality
+In this section, we briefly review the definitions of causality and causal models introduced by HP01b and relevant definitions of causes and explanations in image classification by CH24.
+The reader is referred to Hal19 for further reading.
 We assume that the world is described in terms of variables and their values.
@@
 For simplicity, we assume here that $\mathcal{}V$} is finite, as is ${\mathcal{}R}(Y)$ for every endogenous variable $Y\in{\mathcal{}V}$.
-The set ${\mathcal{}F}$ associates with each endogenous variable $X\in{\mathcal{}V}$ a function denoted $F_{X}$ (i.e., $F_{X}={\mathcal{}F}(X)$) such that $F_{X}:(\times_{U\in{\mathcal{}U}}{\mathcal{}R}(U))\times(\times_{Y\in{\mathcal{}V}-\{X\}}{\mathcal{}R}(Y))\rightarrow{\mathcal{}R}(X)$.
+The set $\mathcal{}F$} associates with each endogenous variable $X\in{\mathcal{}V}$ a function denoted $F_{X}$ (i.e., $F_{X}={\mathcal{}F}(X)$) that expresses the dependency of $X$ on other variables.
+If all variables in $M$ are Boolean, the model is called a binary model.
 The structural equations define what happens in the presence of external interventions.
 Setting the value of some variable $X$ to $x$ in a causal model $M=({\mathcal{}S},{\mathcal{}F})$ results in a new causal model, denoted $M_{X\leftarrow x}$, which is identical to $M$, except that the equation for $X$ in $\mathcal{}F$} is replaced by $X=x$.
-Probabilistic causal models are pairs $(M,\Pr)$, where $M$ is a causal model and $\Pr$ is a probability on the contexts.
-A causal model $M$ is recursive (or acyclic) if its causal graph is acyclic.
-If $M$ is an acyclic causal model, then given a context, that is, a setting $\vec{u}$ for the exogenous variables in ${\mathcal{}U}$, the values of all the other variables are determined.
-In this paper we restrict the discussion to recursive models.
+Some salient features of causal models can be captured by causal graphs, where the nodes represent variables, and the arrows depict the direction of causality.
+A causal model $M$ is recursive if its causal graph is acyclic.
+If $M$ is a recursive causal model, then given a context, that is, a setting $\vec{u}$ for the exogenous variables in $\mathcal{}U$}, the values of all the other variables are determined.
+In this paper, following the convention, we restrict the discussion to recursive models.
 We call a pair $(M,\vec{u})$ consisting of a causal model $M$ and a context $\vec{u}$ a (causal) setting.
@@
 Finally, $(M,\vec{u})\models[\vec{Y}\leftarrow\vec{y}]\varphi$ if $(M_{\vec{Y}=\vec{y}},\vec{u})\models\varphi$, where $M_{\vec{Y}\leftarrow\vec{y}}$ is the causal model that is identical to $M$, except that the variables in $\vec{Y}$ are set to $Y=y$ for each $Y\in\vec{Y}$ and its corresponding value $y\in\vec{y}$.
+The setting of $\vec{Y}$ to $\vec{y}$ effectively cuts the incoming causal transitions into $\vec{Y}$, replacing the variables in $\vec{Y}$ with the values $\vec{y}$.
+These values are then propagated to the variables that causally depend on $\vec{Y}$.
+This change of the values of some of the variables is called an intervention.
 A standard use of causal models is to define actual causation: that is, what it means for some particular event that occurred to cause another particular event.
-There have been a number of definitions of actual causation given for acyclic models (e.g., (?????????)).
-In this paper, we focus on what has become known as the modified Halpern–Pearl definition and some related definitions introduced by (?).
-We briefly review the relevant definitions below.
-The events that can be causes are arbitrary conjunctions of primitive events (formulas of the form $X=x$); the events that can be caused are primitive events, denoting the output of the model.
-Definition 3.1 (Actual cause).
-$\vec{X}=\vec{x}$ is an actual cause of $\varphi$ in $(M,\vec{u})$ if the following three conditions hold:
-AC1.
-$(M,\vec{u})\models(\vec{X}=\vec{x})$ and $(M,\vec{u})\models\varphi$.
-AC2.
-There is a a setting $\vec{x}^{\prime}$ of the variables in $\vec{X}$, a (possibly empty) set $\vec{W}$ of variables in ${\mathcal{}V}-\vec{X}^{\prime}$, and a setting $\vec{w}$ of the variables in $\vec{W}$ such that $(M,\vec{u})\models\vec{W}=\vec{w}$ and $(M,\vec{u})\models[\vec{X}\leftarrow\vec{x}^{\prime},\vec{W}\leftarrow\vec{w}]\neg{\varphi}$, and moreover
-AC3.
-$\vec{X}$ is minimal; there is no strict subset $\vec{X}^{\prime}$ of $\vec{X}$ such that $\vec{X}^{\prime}=\vec{x}^{\prime\prime}$ can replace $\vec{X}=\vec{x}^{\prime}$ in AC2, where $\vec{x}^{\prime\prime}$ is the restriction of $\vec{x}^{\prime}$ to the variables in $\vec{X}^{\prime}$.
-In the special case that $\vec{W}=\emptyset$, we get the but-for definition.
-A variable $x$ in an actual cause $\vec{X}$ is called a part of a cause.
-In what follows, we adopt the convention of Halpern and state that part of a cause is a cause.
-The notion of explanation taken from (?) is relative to a set of contexts.
-Definition 3.2 (Explanation).
-$\vec{X}=\vec{x}$ is an explanation of $\varphi$ relative to a set ${\mathcal{}K}$ of contexts in a causal model $M$ if the following conditions hold:
-EX1a.
-If $\vec{u}\in{\mathcal{}K}$ and $(M,\vec{u})\models(\vec{X}=\vec{x})\wedge\varphi$, then there exists a conjunct $X=x$ of $\vec{X}=\vec{x}$ and a (possibly empty) conjunction $\vec{Y}=\vec{y}$ such that $X=x\wedge\vec{Y}=\vec{y}$ is an actual cause of $\varphi$ in $(M,\vec{u})$.
-EX1b.
-$(M,\vec{u}^{\prime})\models[\vec{X}=\vec{x}]\varphi$ for all contexts $\vec{u}^{\prime}\in{\mathcal{}K}$.
-EX2.
-$\vec{X}$ is minimal; there is no strict subset $\vec{X}^{\prime}$ of $\vec{X}$ such that $\vec{X}^{\prime}=\vec{x}^{\prime}$ satisfies EX1, where $\vec{x}^{\prime}$ is the restriction of $\vec{x}$ to the variables in $\vec{X}^{\prime}$.
-(This is SC4).
-EX3.
-$(M,u)\models\vec{X}=\vec{x}\wedge\varphi$ for some $u\in{\mathcal{}K}$.
-4 Theoretical Foundations
+There have been a number of definitions of actual causation given for acyclic models (e.g.
+(beckers21c; GW07; Hall07; HP01b; Hal19; hitchcock:99; Hitchcock07; Weslake11; Woodward03)).
+In this paper, we are focusing on explainability, which is traditionally defined using some form of actual causation.
+As we argue in the next section, in our setting, the causal models have a simple structure, and hence all the relevant definitions are simplified.
+4. Theoretical Foundations of Our Approach
 We approach the first two research questions theoretically.
-RQ1
+RQ1:
 What is a formal rigorous definition of explanation, suitable for image classification?
-We view an image classifier (e.g.
-a neural network) as a probabilistic causal model.
-Specifically, the endogenous variables are taken to be the set $\vec{V}$ of pixels that the image classifier gets as input, together with an output variable that we call $O$.
-The variable $V_{i}\in\vec{V}$ describes the color and intensity of pixel $i$; its value is determined by the exogenous variables.
-The equation for $O$ determines the output of the model as a function of the pixel values.
-Thus, the causal network has depth 222, with the exogenous variables determining the feature variables, and the feature variables determining the output variable.
-We also assume causal independence between the feature variables $\vec{V}$.
-In our setting, $\vec{V}$ is the set of pixels of the input image, with the values capturing the color and intensity of the pixels.
-Pixel independence is a common assumption in explainability tools.
-This is a non-trivial assumption, and it might seem far-fetched, especially if we consider images capturing real objects: if a group of pixels captures, say, a cat’s ear, then a group of pixels below it should capture a cat’s eye.
-However, we argue that it is, in fact, accurate on images.
-Indeed, consider the “Photobombing” dataset, described in Section 6.1.
+Given an image classifier (e.g., a neural network) $\mathcal{N}$ and an input image $x$, we define a binary causal model $M_{\mathcal{N},x}$ as follows.
+The set ${\mathcal{}V}=\vec{V}\cup\{O\}$ of endogenous variables consists of a set $\vec{V}$ corresponding to the set of pixels $P(x)$ of $x$ and the single output variable $O$.
+Essentially, $\vec{V}$ is a mask, indicating which pixels of $x$ are visible and which are occluded, and the output variable $O$ indicates whether the classification of a partially masked image stays the same as of the original image.
+To keep in line with the formal definitions of structural causal models, the values of $\vec{V}$ are determined by the set of exogenous variables $\mathcal{}U$}.
+Assigning $1$ to a variable $v_{i}\in\vec{V}$ means that the pixel $p_{i}$, corresponding to $v_{i}$, has its original value (taken from $x$).
+Assigning 0 to this variable means that $p_{i}$ is masked – replaced with some predefined masking value.
+The masking operation of $\mathcal{}V$} on $P(x)$ is denoted by ${\mathcal{}V}\odot P(x)$ and is the Hadamard product of these sets viewed as matrices of the same size (corresponding to the input size and shape of $\mathcal{N}$).
+The context $\vec{u}$ that assigns all variables in $\vec{V}$ the value $1$ (i.e., none of the pixels are masked) corresponds to the fully unmasked image $x$.
+The value of $O$ is $1$ iff the output of $\mathcal{N}$ on ${\mathcal{}V}\odot P(x)$, the partially masked image defined by applying $\vec{V}$ to $x$, is $\mathcal{N}(x)$ and is 0 otherwise.
+Clearly, $O=1$ if the image is fully unmasked, that is, $(M_{\mathcal{N},x},\vec{u})\models(O=1)$.
+[…]
+Figure 2.
+A depth-22 binary causal model $M_{\mathcal{N},x}$ for an image $x$ and a classifier $\mathcal{N}$.
+$\vec{v}$ is the vector of values of $\vec{V}$.
+The output $O\in\{0,1\}$ indicates whether the classification of the Hadamard product of $x$ and $\vec{V}$ is the same as the original classification.
+We depict the reasoning process of $M_{\mathcal{N},x}$ in Figure 2, omitting the set of exogenous variables.
+The causal model has depth $2$.
+In what follows, we omit the subscript $\mathcal{N},x$} from the causal model notation if it is clear from the context.
+This construction assumes causal independence between the variables in $\vec{V}$.
+This is common to many approaches to causal and counterfactual explainable AI (USL19; SHG20; PSSBF20; MMTS21; Beckers22; CH24), and is the de facto approach in all black-box explainable AI tools.
+We argue, however, that this is an accurate depiction in the case of images, and not just a convenient approximation.
+Image classification models perform classifications over data, not over what the image represents.
+This data encodes correlations which are due to the underlying data production method being causal, i.e.
+the real world.
+Obscuring or permuting pixels does not lead to a cascading change of other pixels values, as one would expect if pixels were causally connected, instead these permutations disrupt correlations in the data only.
+Consider the “Photobombing” dataset, described in Section 6.1.
 This dataset is comprised of partially obscured images, obtained by overlaying random color patches over Imagenet images.
@@
 Indeed, obscuring a part of the input image either by introducing an artificial object or by positioning a real object in front of the primary subject of the classification does not lead to any change in unobscured pixels.
-For example, obscuring a cat’s ear does not lead to any change in an (unobscured) group of pixels below it, capturing a cat’s eye.
-Thus, pixel independence holds on general images.
-We note that pixels are not concepts: a  [^2] D image is a projection of a scene in the  [^3] D real world on a plane; concepts that are present in the object in a real world can be invisible on this projection, hence the pixel independence.
-We refer the reader to (?) for a more in-depth discussion of causal independence between pixel values in image classification.
-We note that the causal independence assumption is not true in other types of inputs, such as tabular or spectral data.
-For those types of inputs, assuming independence is clearly an approximation and might lead to inaccurate results.
-In this paper, however, we focus on images, where causal independence between pixels holds, and it greatly simplifies explanation extraction, as we show below.
-Moreover, as the causal network is of depth 222, all parents of the output variable $O$ are contained in $\vec{V}$.
-Given these assumptions, the probability on contexts directly corresponds to the probability on seeing various images (which the model presumably learns during training).
-Given an input image $I$, the set of contexts ${\mathcal{}K}$ that we consider for an explanation is the set ${\mathcal{}K}_{I}$ obtained by all partial occlusions of $I$, where an occlusion sets a part $\vec{Y}$ of the image to a predefined masking color $\vec{y}$.
-The probability distribution over ${\mathcal{}K}_{I}$ is assumed to be uniform.
-Under the assumptions above, the following definition is equivalent to Definition 3.2, as shown by (?).
-Definition 4.1 (Explanation for image classification).
-An explanation is a minimal subset of pixels of a given input image that is sufficient for the model $\mathcal{N}$ to classify the image, where “sufficient” is defined as containing only this subset of pixels from the original image, with the other pixels set to the masking color.
-As our goal is ranking of the pixels according to their importance for the classification, we only consider singleton causes, that is, causes consisting of one pixel and its value.
-Moreover, for each pixel we only consider two possible values: its original value in the image and a masking color.
-Definition 4.2 (Singleton cause for image classification).
-For an image $x$ classified by the DNN as $f(x)=o$, a pixel $p_{i}$ of $x$ and its value in $x$ is a cause of $o$ iff there exists a subset $P_{j}$ of pixels of $x$ such that the following conditions hold:
-SC1.
-$p_{i}\not\in P_{j}$;
-SC2.
-changing the color of any subset $P^{\prime}_{j}\subseteq P_{j}$ to the masking color does not change the classification;
-SC3.
-changing the color of $P_{j}$ and the color of $p_{i}$ to the masking color changes the classification.
-We call such $P_{j}$ a witness to the fact that $p_{i}$ is a cause of $x$ being classified as $o$.
-Definition 4.2 is equivalent to a part of a cause according to Definition 3.1 in binary causal models with all variables being independent of each other.
-Moreover, under these restrictions, Definition 4.2 is equivalent to the following definition, which is the original definition of an actual cause by (?).
-Definition 4.3 (Original definition of actual cause).
-$\vec{X}=\vec{x}$ is an actual cause of $\varphi$ in $(M,\vec{u})$ if the following three conditions hold:
-AC1.
-$(M,\vec{u})\models(\vec{X}=\vec{x})$ and $(M,\vec{u})\models\varphi$.
-AC2’.
-There exists a partition $(\vec{Z},\vec{W})$ of $\vec{V}$ with $\vec{X}\subseteq\vec{Z}$ and some setting $(\vec{x}^{\prime},\vec{w}^{\prime})$ of the variables in $(\vec{Z},\vec{W})$ such that if $(M,\vec{u})\models Z=z*$ for $Z\in\vec{Z}$, then
-  1. 1.
-    $(M,\vec{u})\models[\vec{X}\leftarrow\vec{x}^{\prime},\vec{W}\leftarrow\vec{w}^{\prime}]\neg\varphi$.
-In other words, changing the values of $(\vec{X},\vec{W})$ from $(\vec{x},\vec{w})$ to $(\vec{x}^{\prime},\vec{w}^{\prime})$ changes $\varphi$ from true to false;
-  2. 2.
-    $(M,\vec{u})\models[\vec{X}\leftarrow\vec{x},\vec{W}\leftarrow\vec{w}^{\prime},\vec{Z}^{\prime}\leftarrow\vec{z}*]\varphi$ for all subsets $\vec{Z}^{\prime}$ of $\vec{Z}$.
-In other words, setting $\vec{W}$ to $\vec{w}^{\prime}$ should have no effect on $\varphi$ as long as $\vec{X}$ is kept at its current value $\vec{x}$, even if all the variables in an arbitrary subset of $\vec{Z}$ are set to their original values in the context $\vec{u}$.
-AC3.
-$\vec{X}$ is minimal; there is no strict subset $\vec{X}^{\prime}$ of $\vec{X}$ that satisfies AC1 and AC2’.
-The equivalence is based on the following observations.
-The minimality is satisfied due to only considering singletons.
-The set $\vec{Z}$ does not play a role due to the independence between the pixels.
-Removing $\vec{Z}$ and the minimality condition from Definition 4.3 we get Definition 4.2.
-Computing causality in image classification is, therefore, NP-complete, based on the above equivalence and the result by (?), who proved NP-completeness of Definition 4.3 for binary causal models.
-Moreover, following the same reasoning, Definition 4.1 is equivalent to the definition of explanation by (?) for binary causal models, which is DP-complete (?).
-Hence, there is little hope to find an efficient algorithm for computing exact explanations for image classification.
-Finally, the following definition of the degree of responsibility holds for the case of image classifiers and singleton causes as in Definition 4.2.
-Definition 4.4 (Simplified responsibility).
-[…]
-RQ2
+Thus, pixel independence holds on the general data representation of images.
+We refer the reader to CH24 for a more in-depth discussion of causal independence between pixel values in image classification.
+We note that the causal independence assumption may not hold for concept bottleneck models and other models which claim to learn causes and not just correlations (kohconcept).
+For these types of models, assuming independence may result in approximation which might lead to inaccurate results.
+As we discuss below, there is a strong correlation between pixels in a given image, which is not surprising given that images usually capture objects in the real world.
+This does not contradict our claim of the lack of causal connection between the pixels in the image, due to the argument above.
+Given a neural network $\mathcal{N}$ and an input image $x$, let $\vec{u}_{1}$ be the context that assigns $1$ to all variables in $\vec{V}$, and let $\vec{u}_{0}$ be the context that assigns 0 to all these variables.
+We introduce the following definition of explanation.
+Definition 4.1 (Single-Context Explanation).
+A subset $\vec{V}_{exp}$ of $\vec{V}$ is a single-context explanation of a classification $\mathcal{N}(x)$ of an input image $x$ by a classifier $\mathcal{N}$ if the following conditions hold:
+EXIM1.:
+$(M,\vec{u}_{0})\models[\vec{V}_{exp}=1](O=1)$.
+EXIM2.:
+$\vec{V}$ is minimal; there is no strict subset $\vec{V}^{\prime}$ of $\vec{V}$ that satisfies EXIM1, where $\vec{v}^{\prime}$ is the restriction of $\vec{v}$ to the variables in $\vec{V}^{\prime}$.
+When there is no confusion, we call single-context explanation simply an explanation.
+As there is a one-to-one correspondence between the variables in $\vec{V}$ and the pixels of $x$, we also call the subset of pixels $P_{exp}$ of $x$ that corresponds to $\vec{V}_{exp}$ a single-context explanation of $\mathcal{N}(x)$.
+In other words, an explanation is a minimal subset of pixels of a given input image $x$ that is sufficient for the model $\mathcal{N}$ to classify the image, with all other pixels masked.
+Note that we do not assume that the classification of a fully masked input by $\mathcal{N}$ is different from $\mathcal{N}(x)$; if they are equal, the (single) explanation is an empty set.
+Definition 4.1 matches other definitions of explanations in the causality landscape.
+So as not to interrupt the flow for the reader, we define and prove the relevant equivalences in Appendix A.
+The reader can safely ignore this appendix, as it is not directly related to the methods and algorithms presented in this paper.
+We now return to the task at hand, which is computing explanations.
+RQ2:
 Can we compute explanations based on the definition above without opening the black box? What is the complexity of such computation?
-[…]
-Observation 4.5.
-Given an image $x$ and its classification $o$, we can calculate the degree of responsibility of each pixel $p_{i}$ of $x$ by directly applying Definition 4.2, that is, by checking the conditions SC1, SC2, and SC3 for all subsets $P_{j}$ of pixels of $x$ and then choosing a smallest witness subset.
-While there is an underlying Boolean formula that determines the classification $o$ given the values of the pixels of $x$, we do not need to discover this formula in order to calculate the degree of responsibility of each pixel of $x$.
-In the next section we introduce a compositional approach to computing approximate explanations.
-The main idea is first to rank pixels according to their approximate degree of responsibility, and then use a greedy approach to extract (approximate) explanations.
-The algorithm for computing an approximate degree of responsibility is based on the notion of a super-pixel $P_{i}$, which is a subset of pixels of a given image.
-Given an image $x$, we partition it into a small number of superpixels and compute their degree of responsibility for the output of the DNN.
-Then, we only refine those superpixels with a high responsibility (exceeding a predefined threshold).
-The scalability of the approach relies on the following observation, which is shown to be heuristically true in our experiments.
-Observation 4.6.
-The pixels with the highest responsibility for the DNN’s decision are located in super-pixels with the highest responsibility.
+Unfortunately, the precise computation of explanations is intractable—see Appendix A for the proof.
+It is of little surprise, as all the existing definitions of explanations are intractable, except in some very simple special cases.
+There is, therefore, a justification for an approximation algorithm for explanations.
+Section 5, described in the next section, computes approximate explanations.
+It uses an approximate ranking of pixels according to their importance for the classification of the input image as a heuristic to guide explanation discovery.
+We formalize the notion of importance as follows.
+Definition 4.2 (Sufficient responsibility).
+The degree of sufficient responsibility of an explanation $\vec{V}_{exp}$ for the classification of $x$ by $\mathcal{N}$ is defined as $1/|\vec{V}_{exp}|$.
+We also extend this value to all pixels in $\vec{V}_{exp}$.
+That is, $v_{i}$ (and its matching pixel $p_{i}$) have the degree of sufficient responsibility $1/|\vec{V}_{exp}|$ for the classification of $x$ by $\mathcal{N}$, where $\vec{V}_{exp}$ is the smallest explanation for the classification of $x$ that contains $v_{i}$.
+If there is no explanation that contains $v_{i}$, then its degree of sufficient responsibility is defined as 0.
+Similarly to the definition of explanation, Definition 4.2 has a match in the existing causality landscape, as we show in Appendix A, where we also prove its intractability.
+The following observation describes a brute-force approach to computing sufficient responsibility, which is clearly exponential in the number of pixels of the image.
+Observation 4.3.
+Given an image $x$ and its classification $o$, we can calculate the degree of sufficient responsibility of each pixel $p_{i}$ of $x$ by directly applying Definition 4.1, that is, by checking the conditions EXIM1 and EXIM2 for all subsets of pixels of $x$.
+In the next section, we describe an efficient approximation algorithm for computing explanations that is based on an efficiently computable approximate degree of responsibility.
+5. The rex Algorithm
+Refer to captionCausalRankingAlgorithmRefer to captionExplanationextractionRefer to captionRefer to caption
+Figure 3.
+High level overview of Algorithm 3.
+The causal ranking algorithm produces an approximate responsibility map (①).
+The pixels in the image are then ordered by their approximate responsibility, and the explanation extraction algorithm uses this ranking to produce an approximately minimal sufficient explanation (②), which captures the information required for the DNN to give the classification ('starfish' in this example).
+In this section we introduce our algorithm for computing approximate explanations.
+The high-level view of the algorithm is presented in Figure 3.
+The algorithm consists of two independent parts.
+The first part ranks pixels according to their approximate degree of sufficient responsibility (Definition 4.2); the second part is a greedy algorithm that extracts (approximate) explanations from this ranking.
+The algorithm for computing an approximate degree of responsibility is based on the notion of a superpixel $P_{i}$.
+We slightly abuse the definition of a superpixel as used in lime, where it refers to pixels which are clustered via some segmentation algorithm.
+In our context, however, it is simply some subset of pixels of a given image which have not been masked.
+rex does not rely on any externally provided initial segmentation.
+The algorithm instead partitions an input image into a small number of rectilinear superpixels, computes their degree of sufficient responsibility for the output, and then iteratively refines the superpixels with responsibility exceeding some predefined threshold.
+The computation of the degree of sufficient responsibility of each superpixel for the output is precise and is described in Section 5.1.
+The approximation comes from the iterative refinement step in the algorithm, described in Section 5.2, and is due to the fact that the final ranking of pixels depends on the selected partition.
+To ameliorate the effect of a particular partition, the algorithm is repeated a number of times with partitions selected independently at random, and the results are averaged across all the partitions.
+Section 5.3 describes the averaging process and the greedy construction of an approximate sufficient explanation from the ranked list of pixels.
+Section 5.4 gives a step-by-step example to illustrate the working of the algorithm.
+The scalability of the approach relies on the following observation.
+Observation 5.1.
+The pixels with the highest responsibility for the model’s decision are located in superpixels with the highest responsibility.
 Intuitively, the observation holds when pixels with high responsibility do not appear in the superpixels surrounded by other pixels with very low responsibility for the input image classification outcome.
-While this can happen in principle, we do not encounter this case in practice owing to the continuous nature of images (even when the explanation is non-contiguous).
-This property is key to the success of our algorithm.
-5 Compositional Explanations
-Refer to captionCausalRankingAlgorithmRefer to captionExplanationextractionRefer to captionRefer to caption
-Figure 2: High level overview of Algorithm 3.
-Iterating parts 1 and 2 of the algorithm produce a responsibility map (①).
-The pixels in the image are then ordered by their approximate responsibility and used to produce a minimal sufficient explanation (②), which captures all that the DNN needs to give a classification 'starfish'.
-We recall the third research question.
-RQ3
-Is there an efficiently computable approximation to explanations?
-We answer this question affirmatively, by presenting our greedy compositional explanation (CE) algorithm.
-The general idea is to calculate the responsibility of a superpixel and recursively distribute this responsibility to all pixels within this superpixel.
-The CE approach in this work consists of the following steps.
-  1. 1.
-    Given a set of superpixels, compute the responsibility of each superpixel (Section 5.1).
-  2. 2.
-    Following the responsibility result in Step 1, further refine the superpixel and calculate the responsibility for the refined superpixels (Section 5.2).
-We note that the ranking depends on the selected partition.
-To ameliorate the effect of a particular partition, the algorithm is re-run a number of times with partitions selected independently at random, and the results are averaged across all the partitions (see Section 5.3).
-Section 5.4 gives a step-by-step example to illustrate the working of the algorithm.
-5.1 Computing the responsibility of a superpixel
+Although this can happen in principle, especially in the case of adversarial images, we do not encounter this in practice.
+While the model has access to data only, this data has strong correlations due to the underlying data production mechanism.
+5.1. Computing the degree of responsibility of a superpixel
 Given a set of pixels $\mathcal{P}$, we use $\mathds{P}_{i}$ to denote a partition of $\mathcal{P}$, that is, a set $\{P_{i,j}:\bigcup P_{i,j}=\mathcal{P}$ and $\forall j\not=k,P_{i,j}\cap P_{i,k}=\emptyset\}$.
-The number of elements in $\mathds{P}_{i}$ is a parameter, denoted by $s$; in this work, we consider $s=4$.
+The number of elements in $\mathds{P}_{i}$ is a parameter, denoted by $s$; in our implementation, rex, we set $s=4$ as this provides convenient partitioning.
 We refer to $P_{i,j}$ as superpixels.
@@
 Therefore we create combinations of superpixels from the power set of $\mathds{P}_{i}$, denoted $2^{\mathds{P}_{i}}$.
-For an NN $\mathcal{N}$, an input $x$, and a partition $\mathds{P}_{i}$, we can generalize Def.
-4.2 to the set of superpixels defined by $\mathds{P}_{i}$.
-We denote by $r_{i}(P_{i,j},x,\mathcal{N}(x))$ the degree of responsibility of a superpixel $P_{i,j}$ for $\mathcal{N}$’s classification of $x$, given $\mathds{P}_{i}$.
+For an NN $\mathcal{N}$, an input $x$, and a partition $\mathds{P}_{i}$, we can generalize Definition 4.2 to the set of superpixels defined by $\mathds{P}_{i}$.
+We denote by $r_{i}(P_{i,j},x,\mathcal{N}(x))$ the sufficient responsibility of a superpixel $P_{i,j}$ for $\mathcal{N}$’s classification of $x$, given $\mathds{P}_{i}$.
 For a partition $\mathds{P}_{i}$, we denote by $X_{i}$ the set of mutant images obtained from $x$ by masking subsets of $2^{\mathds{P}_{i}}$, and by $\tilde{X}_{i}$ the subset of $X_{i}$ that is classified as the original image $x$.
@@
 $\tilde{X}_{i}=\{x_{m}:\mathcal{N}(x_{m})=\mathcal{N}(x)\}.$
-We compute $r_{i}(P_{i,j},x,\mathcal{N}(x))$, the responsibility of each superpixel $P_{i,j}$ in the classification of $x$, in Algorithm 1. For a superpixel $P_{i,j}$, we define the set
+We compute $r_{i}(P_{i,j},x,\mathcal{N}(x))$, which is an approximation of the degree of sufficient responsibility (Definition 4.2) of each superpixel $P_{i,j}$ for the classification of $x$, in Algorithm 1.
+For a superpixel $P_{i,j}$, we define the set
 $\tilde{X}_{i}^{j}=\{x_{m}:P_{i,j}\mbox{ is not masked in }x_{m}\}\cap\tilde{X}_{i}.$
-[…]
-[…]
+For a mutant image $x_{m}$, we define $\mathit{diff_{i}}(x_{m},x)$ as the number of superpixels in the partition $\mathds{P}_{i}$ that are masked in $x_{m}$ (that is, the difference between $x$ and $x_{m}$ with respect to $\mathds{P}_{i}$).
+For an image $y$, we denote by $y(P_{i,j})$ an image that is obtained by masking the superpixel $P_{i,j}$ in $y$.
+The responsibility of a superpixel $P_{i,j}$ is calculated by Algorithm 1 as a minimum difference between a mutant image and the original image over all mutant images $x_{m}$ that do not mask $P_{i,j}$, are classified the same as the original image $x$, and masking $P_{i,j}$ in $x_{m}$ changes the classification.
+Algorithm 1 $\mathit{responsibility}(x,\mathds{P}_{i})$
 INPUT: an image $x$, a partition $\mathds{P}_{i}$
-OUTPUT: a responsibility map $\mathds{P}_{i}\rightarrow\mathds{Q}$
+OUTPUT: a responsibility map $R:\mathds{P}_{i}\rightarrow\mathds{Q}$
 1:for each $P_{i,j}\in\mathds{P}_{i}$ do
-[…]
+2:  $k\leftarrow\min\limits_{x_{m}}\{\mathit{diff}(x_{m},x)\,|\,\,x_{m}\in\tilde{X}^{j}_{i}\}$
 3:  $r_{i,j}\leftarrow\frac{1}{k+1}$
@@
 5:return $r_{i,0},\dots,r_{i,|P_{i}|-1}$
-5.2 Iterative refinement of responsibility
+5.2. Iterative refinement of responsibility
 Algorithm 1 calculates the responsibility of each superpixel, subject to a given partition.
-Then, it proceeds with only the high-responsibility superpixels.
+It then proceeds with only the high-responsibility superpixels.
+It returns a responsibility map, $R$, which maps superpixels to the rational numbers $\mathds{Q}$, indicating their degree of responsibility.
 Note that in general, it is possible that all superpixels in a given partition have the same responsibility.
@@
 Each quadrant would be equally important for the classification, hence we would not gain any insight into why the image was classified in that particular way.
-In this case, the algorithm chooses another partition.
+In this case, the algorithm starts again from another partition.
 It can also be the case that there might be multiple disjoint combinations of high responsibility superpixels in $\mathds{P}_{i}$.
@@
 Let us say that the passing combination of superpixels $P_{c}=\{0,2\}$, assuming that we have numbered $P_{i,j}$ consecutively, and that $s=4$.
-Each superpixel $P_{i,j}$ has responsibility 0.50.50.5.
-We cannot simply mask superpixel 2 while the refining superpixel 0 as both together are required for the classification.
-We handle this problem by holding one superpixel to its original value while refining the other and then flipping the procedure.
-The superpixel held to its original value, but not being refined, does not get additional responsibility.
-This procedure is relatively scalable to sets containing more than 2 superpixels.
-In our implementation, rex, $s$ is hard-coded to 444 because this result in powerset of superpixels of manageable size and the refinement process for sets of superpixels just discussed needs to handle at most three superpixels at a time.
-[…]
+Each superpixel $P_{i,j}$ has responsibility $0.5$.
+We cannot simply mask superpixel $2$ while refining superpixel 0 as both together are required for the classification.
+We handle this problem by holding one superpixel to its original value while refining the other and then reversing the procedure.
+The superpixel held to its original value, but not being refined, does not get any additional responsibility.
+Algorithm 2 $\mathit{iterative\_responsibility\_refinement}(x,\mathds{P}_{i})$
 INPUT:   an image $x$ and a partition $\mathds{P}_{i}$
-OUTPUT:   a responsibility map $\mathds{P}_{i}\longrightarrow\mathds{Q}$
+OUTPUT:   a responsibility map $R:\mathds{P}_{i}\longrightarrow\mathds{Q}$
 […]
@@
 6:for each $P_{c}\in(2^{\mathds{P}_{i}}-\emptyset)$ s.t $R(P_{c})\not=0$ do
-[…]
+7:  $R^{\prime}\leftarrow R^{\prime}\,\cup\,\mathit{iterative\_responsibility\_refinement}(x,P_{c})$
 8:end for
 9:return $R^{\prime}$
-One partition is rarely sufficient for a high-quality explanation, therefore we compute Algorithm 1 many times and compose the results, as shown in Algorithm 2.
+One partition is rarely sufficient for a high-quality (i.e.
+small) explanation, therefore we compute Algorithm 1 many times and compose the results, as shown in Algorithm 2.
 This calculates the responsibility for each superpixel (Line 1).
@@
 The algorithm terminates when: 1) the superpixels in $\mathds{P}_{i}$ are sufficiently refined (containing only very few pixels), or 2) when all superpixels in $\mathds{P}_{i}$ have the same responsibility (this condition is for efficiency).
-5.3 Explanation extraction
+In particular, if no further subdivision of a superpixel results in the desired classification, responsibility for all superpixels is 0 and the algorithm terminates.
+5.3. Explanation extraction
 So far, we assume one particular partition $\mathds{P}_{i}$, which Algorithm 2 recursively refines and calculates the corresponding responsibilities of superpixels in each step by calling Algorithm 1.
 We note that the choice of the initial partition over the image can affect the values calculated by Algorithm 2, as this partition determines the set of possible mutants in Algorithm 1.
-We ameliorate the influence of the choice of any particular partition by iterating the algorithm over a set of partitions.
-Twenty iterations of the algorithm will therefore yield 202020 starting partitions, chosen at random.
+We ameliorate the influence of the choice of any particular partition by iterating the algorithm over a set of initial partitions.
+Twenty iterations of the algorithm will therefore yield $20$ starting partitions, chosen at random.
 rex allows the user to choose from a number of different types of random to build these partitions, with uniform being the default.
@@
 $\mathds{P}^{x}$.
-[…]
+Algorithm 3 $\mathit{explanation}(x)$
 INPUT:   an input image $x$, a parameter $N\in\mathds{N}$
@@
 […]
-2:for $c$ in 111 to $N$ do
+2:for $c$ in $1$ to NN do
 3:  $\mathds{P}^{x}\leftarrow$ sample a partition
-[…]
+4:  $R\leftarrow\mathit{iterative\_responsibility\_refinement}(x,\mathds{P}^{x})$
 5:  for each $P_{i,j}\in~\mbox{domain of}~R$ do
@@
 […]
-$\mathcal{E}\leftarrow\emptyset$
 […]
+11:for each pixel $p_{i}\in\mathit{pixel\_ranking}$ do
 12:  $\mathcal{E}\leftarrow\mathcal{E}\cup\{p_{i}\}$
-[…]
-[…]
+13:  $x^{\mathit{exp}}\leftarrow$ mask pixels of $x$ that are not in $\mathcal{E}$
+14:  if $\mathcal{N}(x^{\mathit{exp}})=\mathcal{N}(x)$ then
 15:   return $\mathcal{E}$
@@
 After $N$ iterations, all pixels are ranked according to their responsibility $r_{p}$.
-The remainder of Algorithm 3 follows the method for explaining the result of an image classifier by (?).
+The remainder of Algorithm 3 follows the method for explaining the result of an image classifier by sun2020explaining.
 That is, we construct a subset of pixels $\mathcal{E}$ to explain $\mathcal{N}$’s output on this particular input $x$ greedily.
@@
 In practice, while our algorithm uses an iterative average of a greedy approximation, it yields highly accurate results (Section 6).
-Furthermore, our approach is simple and general, and uses the neural network as a blackbox.
-5.4 Illustrative example
+Furthermore, our approach is simple and general, and uses the neural network as a black-box.
+5.4. Illustrative example
 Refer to captionRefer to captionRefer to caption① initialpartitionsRefer to captionRefer to captionRefer to caption② refinedpartitionsRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption③ responsibilitymapRefer to caption④ extractedexplanation
-Figure 3: The rex algorithm in action: rex creates an initial random partition of an image into 444 sections (①).
-All combinations of these sections are queried by the model, with further random refinement applied to those sections or combinations of sections that meet the requirements (②).
+Figure 4.
+The rex algorithm in action: rex creates an initial random partition of an image into 44 sections (①).
+All combinations of these sections are queried by the model, with further refinement applied to those sections or combinations of sections that meet the requirements (②).
 Some combinations, highlighted in green, are classified as bus, others, in red, are not.
 This results, after several iterations, in a detailed responsibility map (③), from which a minimal passing explanations can be extracted (④).
-To illustrate how rex works, consider Figure 3, which is classified as 'bus' by ResNet50, even though there is an occlusion in the middle.
-Initially, rex picks up an arbitrary partition of the image with four superpixels.
-This results in 151515 combinations of masking superpixels, though in practice, we only need to examine 141414 combinations as the $15^{th}$ is the entire image, which we have already tested.
+To illustrate how rex works, consider Figure 4, which is classified as 'bus' by a ResNet50, even though there is an occlusion in the middle.
+Initially, rex picks an arbitrary partition of the image.
+This results in $15$ combinations of masking superpixels, though in practice, we only need to examine $14$ combinations as the $15^{th}$ is the entire image, which we have already tested.
 We use Algorithm 1 to calculate the responsibility of each superpixel.
-Those superpixels or combinations of superpixels which contribute towards the correct classification of further refined.
+Those superpixels or combinations of superpixels which contribute towards the correct classification are further refined.
 This iterative refinement is a recursive application of the same initial random partitioning into four superpixels, as in Algorithm 2.
-The responsibility mapss shown in Figure 4 demonstrate the importance of Algorithm 3.
+The responsibility maps shown in Figure 5 demonstrate the importance of Algorithm 3.
 The initial partitioning of the image can greatly affect the quality of the responsibility map.
 Figure 5(a) shows the responsibility map after one iteration of the algorithm.
-While it still indicates well the areas of interest in the image it is also quite rectilinear.
-Figure 4(b) contains a lot more information about the distribution of responsibility over the pixels in the image.
+While it still indicates well the areas of interest in the image it is also rather crude.
+Figure 5(b) contains more refined information about the distribution of responsibility over the pixels in the image.
 We now have two distinct peaks of responsibility on either side of the central occlusion (a person).
-Further iterations cause further refinement.
-By iteration 100 (Figure 4(e)), it is clear that the dominant explanation is around the rear tyre of the bus.
-The other explanation however, while depressed in relation to the tyre, has not vanished and can still be found through searching the responsibility map (?).
-Indeed, a disjoint explanation exists for this image, as given in Figure 5(a).
+Further iterations bring about further refinement.
+By iteration $100$ (Figure 5(e)), it is clear that the ‘dominant’ explanation is around the rear tire of the bus.
+The other explanation however, while depressed in relation to the tire, has not vanished and can still be found through searching the responsibility map (CKK25).
+Moreover, a disjoint explanation exists for this image, as given in Figure 6(a).
 (a) $N=1$
@@
 (e) $N=100$
-Figure 4: Improvement of rex’s pixel ranking as the number of iterations $N$ increases (Algorithm 3)
+Figure 5. Improvement of rex’s pixel ranking on ‘bus’ (Figure 4) as the number of iterations NN increases (Algorithm 3)
 (a) Non-contiguous explanation
 (b) Contiguous explanation
-Figure 5: Two different explanations extracted from the 303030 iteration ranking in Figure 4Figure 5(a) contrasts interestingly with Figure 5(b): it seems that the front tire by itself is not sufficient (a very small rectangular section of the rear tire is included in the explanation), but the rear tire is.
-5.5 Termination and complexity
-It is clear that the algorithm terminates as soon as the parts can no longer be divided into superpixels.
+Figure 6.
+Two different explanations extracted from the $30$ iteration ranking in Figure 5Figure 6(a) contrasts interestingly with Figure 6(b): it seems that the front tire by itself is not sufficient (a very small rectangular section of the rear tire is included in the explanation), but the rear tire alone is sufficienct.
+5.5. Termination and complexity
+It is clear that the algorithm necessarily terminates as soon as the parts can no longer be divided into superpixels.
 Moreover, the number of calls it performs to the model is linear in the size of the image, as proved in the following lemma.
@@
 The number of steps is determined by the termination condition on the size of a single superpixel, which in the worst case is the same as a single pixel.
-In our setting, the algorithm terminates when a single superpixel is smaller contains fewer than 101010 pixels111rex also terminates when no combination of active superpixels has the required classification or a user-defined work budget is exceeded..
+In our setting, the algorithm terminates when a single superpixel is contains fewer than $10$ pixels111rex also terminates when a user-defined work budget is exceeded..
 However, in general, the algorithm can continue down to the level of a single pixel, thus resulting in $n$ pixels in the last step.
@@
 ∎
-6 Evaluation
+Recall the research question RQ3:
+RQ3:
+Is there an efficiently computable approximation to explanations?
+The complexity analysis of Algorithm 3 in Lemma 5.2 provides an affirmative answer to this research question.
+6. Evaluation
 We conduct a large scale investigation into rex, comparing it with a host of popular XAI tools.
@@
 Just as there is no universally agreed definition of explanation, there is also no single best way to evaluate the quality of an explanation.
-RQ4
-What are suitable quality measures for explanations? A measure is suitable if it is meaningful with respect to quality and produces results that match our intuition.
-A causal explanation is a minimal set of pixels from an image which are sufficient to obtain the same top classification as the complete set of all pixels of an image.
+RQ4:
+What are suitable quality measures for explanations?
+A causal explanation is a minimal set of pixels from an image which are sufficient to obtain the same top-$1$ classification as the complete set of all pixels of the image.
 This immediately suggests size of explanation as a robust quality measure.
-As none of the tools (apart from rex) compute minimal, sufficient explanations, we use the mechanism of extraction (Section 5.3) on the output of all tools tested.
-We also use the same approach to compute insertion and deletion curves over all XAI outputs.
-Utility and interpretation of deletion curves
-(a) Starfish
-(b) Bison
-Figure 6: An image with a high deletion curves does not indicate a low quality explanation.
-The starfish (Figure 6(a)) has a normalized deletion of 0.320.320.32 because their are multiple points of high responsibility in the image.
-These points indicate the presence of multiple explanations.
-The bison (Figure 6(b)) has a low deletion of 0.10.10.1 because there is only one point of interest in the image.
-Deletion is a difficult measure to interpret.
-When insertion and deletion curves were introduced in rise (?), the authors stated that a high insertion and a low deletion is preferable.
-This assertion relies on the assumption that there is only one explanation for the image’s classification, or that the XAI tool finds only one explanation.
-Moreover, saliency values outside of the single ‘explanation’ must be strictly uninformative: in order to keep deletion low, the saliency map cannot provide information about any substructures present in the image.
-Pixels must be added back into the image in an order that is essentially random.
-rex does not work like this: the responsibility map provides detail over the entire pixel space.
-As an example, if we examine Figure 6(a), we have an image with a normalized insertion curve of 0.970.970.97 but a normalized deletion curve of 0.320.320.32.
-There would appear to be tension between the two numbers, as a high insertion suggests a high quality explanation whereas a high deletion suggests a low quality explanation.
-The mystery is resolved, however, by noting that images can and do contain multiple explanations, as first shown in (?).
-Indeed, there are four distinct peaks of responsibility in Figure 6(a), each of which corresponds to an independent explanation for starfish (Figure 7).
-(a) Explanation 1
-(b) Explanation 2
-(c) Explanation 3
-(d) Explanation 4
-Figure 7: The responsibility map rex provides is sufficiently detailed to extract multiple valid explanations for most images. These starfish explanations were extracted from Figure 6(a)).
-The deletion score is high because the responsibility map is informative even for the more lowly ranked pixels.
-In the multiple explanation scenario, one would expect a high deletion score.
-Uniquely among the tools examined here, rex supports multiple explanations for an image natively (?).
-In Tables 1, 3 and 5 we follow the usual procedure of indicating the lowest deletion curve value, but stress that lowest does not necessarily indicate best.
-We quantify the quality of the explanations with several complementary methods.
-First, we calculate the minimal explanation size as a percentage of the size of the image.
-This coincides with Definition 3.2.
-It also gives an idea of how concentrated the responsibility or saliency maps provided by the XAI tools are.
-The smaller the area of the explanation, the more the XAI tool was able to isolate precisely those pixels (and only those pixels) required to match the original classification.
-Furthermore, we calculate normalized insertion and deletion curves for all images.
-As stated above, it is not clear how to compare the area under curve (AUC) of a low confidence image against a high confidence image to derive an average AUC performance.
-It is, after all, not the fault of the XAI tool that the confidence of the model is low on any particular image.
-To overcome this problem, we follow (?) and normalize all curves by the initial confidence of the model on image under test.
-Note that this can result in the the insertion curve AUC being greater than 111.
-This phenomenon occurs when the confidence on the overall image is lower than the confidence of intermediate insertion stages.
+As none of the tools (apart from rex) compute minimal, sufficient explanations, we use our mechanism of extraction (Section 5.3) on the output of all tools tested.
+We also use a number of complementary measures as proxies for the quality of the explanations.
+Insertion and deletion curves, introduced in petsiuk2018rise, are widely used to assess the quality of saliency maps.
+For insertion curves, the model confidence on a target class is measured as pixels are inserted over a baseline value.
+The order of insertion is derived from the map.
+If a map has accurately identified the most important pixels for a class, then the class confidence should rise quickly.
+This results in a large AUC (area under curve).
+Deletion curves are calculated in the opposite way, replacing the pixels with a baseline value.
+Again, if the map has identified the most important pixels accurately, then the model confidence should drop quickly, resulting in a low AUC.
+While a rex explanation is not a map, we calculate insertion and deletion over the responsibility map.
+The AUC for both insertion and deletion are obviously tied to the original confidence of the model.
+To allow for a fair comparison between images with different initial confidences, we follow calderon2024real and normalize all curves by the initial confidence of the model on the image under test.
+Note that this can result in the the insertion curve AUC being greater than $1$.
+This phenomenon occurs when the confidence on the entire image is lower than the confidence of intermediate insertion stages.
 This makes intuitive sense: if the pixel ranking is accurate, then those pixels that either do not contribute towards the classification, or reduce confidence, are left to be added last.
-Without these pixels, confidence is potentially much higher.
+Without these ‘negative’ pixels, confidence is potentially much higher.
 Comparison with human segmentation
-For those datasets that contain segmentation information, we also measure the intersection of the minimal explanation with the segmentation.
+For those datasets that contain segmentation information, we also measure the intersection of the minimal explanation with that segmentation.
 This segmentation is human-provided and we argue corresponds most closely with what a human considers acceptable.
-On VOC2012, for example, we would like an minimal explanation to reside mostly inside the human-provided segmentation mask and therefore contain only a small number of pixels outside the mask.
-For datasets that feature occluded objects, a good explanation would have as few pixels from inside the occlusion as possible, as the occlusion should not make a contribution towards the model’s classification.
-To answer RQ4, we argue that the size of explanations, as well as insertion curves and the intersection with the occlusion are measures that match our intuition.
-However, we argue that deletion curves are not, as they remain high in presence of multiple explanations.
-6.1 Experimental Setup
+On VOC2012, for example, we would like a minimal explanation to reside mostly inside the human-provided segmentation mask and therefore contain only a small number of pixels outside the mask.
+For datasets that feature occluded objects, a good explanation would have as few pixels from inside the occlusion region as possible, as the occlusion should not make a substantial contribution towards the model’s classification.
+To answer RQ4, we argue that the size of explanations, as well as insertion curves and the intersection with the human segmentation are measures that match our intuition.
+However, we argue that deletion curves are not, as they remain relatively high in presence of multiple detected explanations (Section 6.3).
+6.1. Experimental Setup
 We answer research questions RQ5–RQ7 experimentally.
 We have implemented the proposed explanation approach in the publicly available tool rex.
-For all other XAI tools, apart from rise, we used the Captum PyTorch library (?) [^2] https://www.captum.ai.
-For rise, we used the authors’ implementation [^3] https://github.com/eclique/RISE, which we lightly altered to use PyTorch rather than tensorflow.
-In the evaluation, we compare rex with a wide range of explanation tools, specifically Grad-CAM (?), KernelShap (?), GradientShap (?), rise (?), lime (?), ig (?), noisetunnel (?).
-For the ResNet50 model only, we also include lrp.
-We were unable to make the captum version of lrp run on either the ViT or the ConvNext model.
+For all other XAI tools, apart from rise, we used the Captum PyTorch library (kokhlikyan2020captum) [^2] https://www.captum.ai.
+For rise, we used the authors’ implementation [^3] https://github.com/eclique/RISE, which we lightly altered to use PyTorch rather than the original TensorFlow framework.
+In the evaluation, we compare rex with a wide range of explanation tools, specifically Grad-CAM (CAM), KernelShap (lundberg2017unified), GradientShap (gradientshap), rise (petsiuk2018rise), lime (lime), ig (sundararajan2017axiomatic) and noisetunnel (smilkov2017smoothgradremovingnoiseadding).
+For the ResNet50 model only, we also include lrp (LRP).
+We were unable to make the captum version of lrp run on the other tested models.
 While our main interest is in black-box explainability, we have included methods which require access to the model gradient, such as GradientShap and ig, and internal model layers, i.e.
 Grad-CAM.
-We call those tools which require internal access to the model ‘white-box’; those that require the access to the gradient we term ‘grey-box’, and those that just need the model’s classification, we call ‘black-box’.
-rex is a black-box method.
-PyTorch permits optional access to the gradient, so tools such as ig require at least some flexibility when calling the model.
-This is not guaranteed with proprietary models: hence our grey-box category.
-We use 444 data sets: ImageNet-1k-mini validation (?), ECSSD (?), Pascal VOC2012 (?), and a “Photobombing” dataset we created.
+This is largely due to the relative paucity of purely black-box XAI methods for image data.
+We use $4$ data sets: ImageNet-1k-mini validation (imagenet), ECSSD (ecssd), Pascal VOC2012 (pascal-voc-2012), and a “Photobombing” dataset we created.
 Imagenet1k-mini comes with labels for ground truth.
@@
 All experiments were conducted using a server running Ubuntu 20.04 with an Nvidia A40 GPU.
-Moreover, all tools were used with default settings.
+All tools were used with default settings.
 For rex in particular, this means that it performs $20$ iterations of the algorithm ($N$ in Algorithm 3), with a minimum superpixel size of $10$ pixels.
-The pruning strategy for the work queue is “area” (the passing mutants are ordered by size) and only one item is kept in the work queue at a time, so we only refine the smallest passing mutant.
+The pruning strategy for the work queue is “area” (the passing mutants are ordered by size, smallest to largest) and only one item is kept in the work queue at a time, so we only refine the smallest passing mutant.
 Partitions are created uniformly at random.
@@
 This is practical for images as we can split the image with just one call to a random number generator.
-The number of mutants produced is also relatively small (151515), which can usually be fit into a single batch for model processing.
-Having more initial superpixels does not lead to greater expressivity: we iteratively refine passing combinations of superpixels further, essentially recreating a more detailed initial partition produced by a greater number of superpixels.
-The termination conditions for the partition refinement in Algorithm 2 are: 1) the area of a superpixel is less than 101010 pixels of (resized) input image, 2) the four superpixels share the same responsibility or 3) no combination of superpixels produces the desired classification.
-6.2 Results
-RQ5
+The number of mutants produced is also relatively small ($15$), which can usually be fit into a single batch for model inference.
+Having more initial superpixels does not lead to greater expressivity: we iteratively refine passing combinations of superpixels, essentially recreating the more detailed initial partition which would be produced by a greater number of superpixels.
+The termination conditions for the partition refinement in Algorithm 2 are: 1) the area of a superpixel is less than $10$ pixels of (resized) input image, or 2) the four superpixels share the same responsibility.
+6.2. Results
+RQ5:
 What is the precision of explanations computed by our algorithm compared to other XAI methods?
-RQ6
+RQ6:
 Is there a trade-off between precision and compute cost of the explanations?
-RQ7
+RQ7:
 Can black-box methods achieve the same quality of explanations as white and grey-box methods?
-Table 1: Results for ResNet50 model.
+Table 1. Results for ResNet50 model.
 (a) Results on Voc
@@
 rise 0.1165 0.1599 0.95 0.2752 0.5451
-Tables 1, 3 and 5 show the experimental results of evaluation of rex against 888 other XAI tools over 444 datasets, with three different models.
-The columns in the tables are: mean area (area), $\sigma$ of area (standard deviation), normalized insertion curves AUC (ins), normalized deletion curves AUC (del), proportion of explanation inside the relevant segment (IN), and the fraction of the explanation that is outside the relevant segment (OUT), shown only where appropriate.
+Tables 1, 3 and 5 show the experimental results of evaluation of rex against $8$ other XAI tools over $4$ datasets, with three different models.
+The columns in the tables are: mean area (area), $\sigma$ of area (standard deviation), normalized insertion curve AUC (ins), normalized deletion curve AUC (del), proportion of explanation inside the relevant segment (IN), and the fraction of the explanation that is outside the relevant segment (OUT), shown only where appropriate.
 Bold indicates best result in category for all columns.
-Note that while we have indicated lowest AUC for deletion, lower here is not necessarily better, as we argued earlier.
+Note that while we have indicated lowest AUC for deletion as ‘best’, lower here is not necessarily better, as we argue in Section 6.3.
 Resnet
 Table 1 shows the results with the ResNet50 model.
-rex consistently produces the most precise explanations.
+rex consistently produces the most smallest explanations.
 This is reinforced by the observations that rex also produces the highest insertion AUC of any of the tools.
-rex does particularly well on the ‘Photobombing‘ dataset, where its explanations are almost entirely (0.970.970.97) disjoint with the inserted occlusion (Figure 12).
+rex does particularly well on the ‘Photobombing‘ dataset, where its explanations are almost entirely ($0.97$) disjoint from the inserted occlusion (Figure 13).
 Figure 7 shows that rex has the most concentrated explanations and consistent results, with very few outliers.
 The grey-box tools perform worse than the other methods.
-Note that rex even outperforms Grad-CAM.
-Table 3: Results for ViT model.
+Note that rex even outperforms Grad-CAM, a white-box method.
+Table 3. Results for ViT model.
 (a) Results on Voc
@@
 ViT
-Table 3 shows the results with the ‘vit_b_32‘ model from torchvision.
-Explanations across all tools are much larger than for ResNet50model, which seems to be a feature of the model.
+Table 3 shows the results with the ‘vit_b_32’ model from torchvision.
+Explanations across all tools are much larger than for the ResNet50, which seems to be a feature of this model.
 rex, however, still manages to produce explanations with the lowest number of redundant pixels.
 In general, rex, lime and Grad-CAM are the best performing tools.
-Even in cases where lime slightly outperforms rex (e.g., on ECSSD IN), with 0.730.730.73 of its explanation inside the segmentation against 0.720.720.72, the effect is very small.
+Even in cases where lime slightly outperforms rex (e.g., on ECSSD IN), with $0.73$ of its explanation inside the segmentation against $0.72$, the effect is very small.
 Figure 8 shows that rex again has the most concentrated explanations and consistent results, with very few outliers.
@@
 The significant difference in explanation size between the ResNet model and ViT model is, perhaps, of independent interest.
-Table 5: Results for convnext model.
+Table 5. Results for convnext model.
 (a) Results on Voc
@@
 ConvNext
-Table 5 shows the results with the ‘convnext_large‘ model from torchvision.
+Table 5 shows the results with the ‘convnext_large’ model from torchvision.
 rex, lime, Grad-CAM and rise are the best performing tools.
-It is interesting to note that 333 of these tools fall into our ‘black-box‘ category.
+It is interesting to note that $3$ of these tools fall into our black-box category.
 Figure 9 shows that rex has the best performance again, this time slightly ahead of Grad-CAM, as Grad-CAM has many more outliers than rex and a slightly larger inter-quartile range.
-Explanation size is comparable with that of the ResNet model, making the ViT model somewhat of an outlier.
-Figure 8: Box plot of all tools over all datasets with a ResNet50 model. rex has the lowest median value and also the smallest number of outliers.
-Figure 9: Box plot of all tools over all datasets with a ViT model.
+Explanation size is comparable with that of the ResNet model, revealing the ViT model to be somewhat of an outlier.
+Figure 7. Box plot of all tools over all datasets with a ResNet50 model. rex has the lowest median value and also the smallest number of outliers.
+Figure 8.
+Box plot of all tools over all datasets with a ViT model.
 rex has the lowest median value and also the smallest number of outliers.
 ViT explanations are noticeably larger than those for ResNet50 and ConvNext.
-Figure 10: Box plot of all tools over all datasets with a ConvNext model.
+Figure 9.
+Box plot of all tools over all datasets with a ConvNext model.
 rex has the lowest median value and also the smallest number of outliers.
@@
 As all tools were evaluated on the same hardware with the same models and datasets, the results should be consistent.
-The white and gray-box methods are clearly the fastest, as expected.
+The white and grey-box methods are clearly the fastest, as expected.
 Indeed, Grad-CAM is essentially instantaneous, as the overhead over one call to the model is insignificant.
 No black-box tool can compete with this, by the nature of black-box tools.
-All the gradient based methods took 3$\approx$ 3≈ 3 seconds to produce initial maps.
-While they are still fast, the quality of their explanations is lower.
-Both rex and lime took 4$\approx$ 4≈ 4s to produce maps, with rise being much slower (10$\approx$ 10≈ 10s).
-If we take into account the extra time for the explanation extraction (which only rex does by default), the efficiency picture changes slightly.
-rex’s total compute time increases only slightly (to 6$\approx$ 6≈ 6s) due the the uniformly small size of its explanations.
-Other tools suffer more though, even Grad-CAM’s average compute time increases to 4$\approx$ 4≈ 4s.
+All the gradient-based methods took $\approx 3$ seconds to produce initial maps.
+Note that this is the time taken to generate initial saliency, minimal pixel subsets still need to be extracted.
+Both rex and lime took $\approx 4$s to produce maps, with rise being much slower ($\approx 10$s).
+Taking into account the extra time for the explanation extraction (which only rex does by default), the efficiency picture changes slightly.
+rex’s total compute time increases only slightly (to $\approx 6$s) due the the uniformly small size of its explanations.
+Other tools suffer more, even Grad-CAM’s average compute time increases to a total of $\approx 4$s.
+This is due to the relatively noisy output of Grad-CAM, requiring more work from the extraction algorithm as a result.
 rex is an efficient black-box tool: for RQ6 therefore, our results show that there is little trade-off between quality and compute cost.
-6.3 Discussion
+6.3. Discussion
+(a) Starfish
+(b) Bison
+Figure 10.
+An image with a high deletion curves does not indicate a low quality explanation.
+The starfish (Figure 10(a)) has a normalized deletion of $0.32$ because their are multiple points of high responsibility in the image.
+These points indicate the presence of multiple explanations.
+The bison (Figure 10(b)) has a low deletion of $0.1$ because there is only one point of interest in the image.
+Deletion is a difficult measure to interpret.
+Utility and interpretation of deletion curves
+petsiuk2018rise stated that a high insertion and a low deletion is preferable.
+This assertion relies on the assumption that there is only one explanation for the image’s classification, or that the XAI tool finds only one explanation.
+For deletion AUC to remain low, saliency values outside of the single ‘explanation’ must be strictly uninformative: the saliency map cannot provide information about any other substructures present in the image.
+It must destroy that signal.
+Essentially, pixel deletion after the first discovered explanatory pixel set must be close to random, breaking up information from other potential explanations.
+Given that a large body of work is dedicated to ‘cleaning up’ saliency maps, especially from Grad-CAM-style tools (noisetunnel; smilkov2017smoothgradremovingnoiseadding; gradcamplusplus), it is not surprising that this genuine signal is lost.
+rex does not work like this: the responsibility map provides detailed information over the entire pixel space.
+Because the map is not itself the explanation, there is no motivation to remove or delete ‘inconvenient’ hot-spots.
+Moreover, because rex is not using activation, which can be noisy, every hot-spot in the rex map is indeed associated to some degree with the desired classification.
+rex utilizes the richness of its responsibility map to discover multiple, different explanations of an image (CKK25).
+This is unlike the other tools in our experimental comparison, all of which return one explanation by design.
+As an example, if we examine Figure 10(a), we have an image with a normalized insertion curve of $0.97$ but a normalized deletion curve of $0.32$.
+There would appear to be tension between the two numbers, as a high insertion suggests a high quality explanation whereas a high deletion suggests a low quality explanation.
+The mystery is resolved, however, by noting that the image contains multiple sufficient explanations.
+Indeed, there are four distinct peaks of responsibility in Figure 10(a), each of which corresponds to an independent explanation for starfish (Figure 11).
+It takes a long time to delete all of the pixels from the well-defined peaks in the responsibility map.
+(a) Explanation 1
+(b) Explanation 2
+(c) Explanation 3
+(d) Explanation 4
+Figure 11. The responsibility map rex provides is sufficiently detailed to extract multiple valid explanations for most images. These starfish explanations were extracted from Figure 10(a).
+In Tables 1, 3 and 5 we follow the usual procedure of indicating the lowest deletion curve value, but stress that lowest does not necessarily indicate best.
+Differences between methods
 Broadly speaking, the tools can be split into three different sets.
@@
 (c) ConvNext
-Figure 11: Cumulative plots of explanations sizes with 333 different models over all datasets.
+Figure 12.
+Cumulative plots of explanations sizes with $3$ different models over all datasets.
 rex consistently produces smaller explanations, meaning it identifies precisely those pixels required for the classification, with fewer unnecessary pixels.
@@
 rex does not require such existing segmentation masks.
-(?) conducted an investigation on brain MRI explainability using a different model and different data set and a similar, though not overlapping, set of explainability tools.
+blake2024explainable conducted an investigation on brain MRI explainability using a different model and different data set and a similar, though not overlapping, set of explainability tools.
 On this dataset, lime’s performance is hampered by medically meaningless segmentation.
@@
 This holds because other tools (such as lime) may have similar IN values, but a larger explanation size.
-Figure 12: Photo Bombing images and output from rex
-We evaluated the results produced by rex with those produced by other explainability tools on a number of standard measures proposed in the literature, such as the size of explanations, insertion and deletion curves.
-For the datasets containing the ground truth or some part of it, such as the segmentation information or the partial occlusion, we also measured the percentage of the explanation that is clearly outside of an area where it should be contained (such as the occlusion in partially occluded areas).
+Figure 13. Photo Bombing images and responsibility maps from rex
+We evaluated the results produced by rex against those produced by other explainability tools on a number of standard measures proposed in the literature, such as the size of explanations, insertion and deletion curves.
+For the datasets containing the ground truth or some part of it, such as the segmentation information or the partial occlusion, we also measured the percentage of the explanation that is clearly outside of an area where it should be contained, such as the occlusion in partially occluded areas (Figure 13).
 For all other measures, rex shows superior performance, which answers RQ5.
 Finally, for RQ7, we did not exhaustively compare the results against all possible explainability methods. Rather, we chose the state-of-the-art primary attribution methods from the Captum library.
-7 Conclusions and Future Work
+7. Conclusions and Future Work
 In this paper we described an approach to explainability of image classifiers that is rooted in actual causality and computes explanations based on the formal definitions of sufficient explanations.
-As the exact computation is intractable, we described a modular algorithm for computing approximate explanations using a causal ranking of parts of the input image, viewing the classifier as a black box.
+As the exact computation is intractable, we described a modular algorithm for computing approximate explanations using a causal ranking of parts of the input image, viewing the classifier as black-box.
 Our experiments demonstrate that an implementation of our algorithm in the tool rex produces results that are superior to those of state-of-the-art tools according to standard measures.
-In the future, we will apply rex to other modalities that lend themselves to occlusion-based reasoning, such as tabular data, time series, spectroscopy, etc.
-We will also explore different domains of application, where precise explanations are instrumental, in particular mission-critical and safety-critical domains, such as the healthcare AI and autonomous vehicles.
+In the future, we will apply rex to other modalities that lend themselves to occlusion-based reasoning, such as tabular data, time series, and spectroscopy.
+We will also explore different domains of application, where precise explanations are instrumental, in particular mission-critical and safety-critical domains, such as in healthcare AI and autonomous vehicles.
 The healthcare domain will require some adaptations of our algorithms, since the quality of medical images is different from the quality of general images (e.g., there are fewer colors in an X-ray than in a general image).
 In the automotive domain, we will adapt rex to analyze object detectors and also devise new algorithms that work in (near) real-time, to address the needs of explainability for autonomous vehicles.
+Appendix A Our Definitions in the Actual Causality Landscape
+For the ease of presentation, we first restate our definition of explanation. See 4.1
+The following is the latest definition of actual cause due to Halpern (Hal15) (see (HP01b) for the original definition).
+Definition A.1.
+[Actual cause (Hal15)] $\vec{X}=\vec{x}$ is an actual cause of $\varphi$ in $(M,\vec{u})$ if the following three conditions hold:
+AC1.:
+$(M,\vec{u})\models(\vec{X}=\vec{x})$ and $(M,\vec{u})\models\varphi$.
+AC2.:
+There is a a setting $\vec{x}^{\prime}$ of the variables in $\vec{X}$, a (possibly empty) set $\vec{W}$ of variables in ${\mathcal{}V}-\vec{X}^{\prime}$, and a setting $\vec{w}$ of the variables in $\vec{W}$ such that $(M,\vec{u})\models\vec{W}=\vec{w}$ and $(M,\vec{u})\models[\vec{X}\leftarrow\vec{x}^{\prime},\vec{W}\leftarrow\vec{w}]\neg{\varphi}$, and moreover
+AC3.:
+$\vec{X}$ is minimal; there is no strict subset $\vec{X}^{\prime}$ of $\vec{X}$ such that $\vec{X}^{\prime}=\vec{x}^{\prime\prime}$ can replace $\vec{X}=\vec{x}^{\prime}$ in AC2, where $\vec{x}^{\prime\prime}$ is the restriction of $\vec{x}^{\prime}$ to the variables in $\vec{X}^{\prime}$.
+The following lemma states their equivalence in our setting.
+Lemma A.2.
+Given a image classifier $\mathcal{N}$, an input $x$, and a causal model $M$ derived from them as described in Section 4, a subset $\vec{V}_{exp}$ of $\vec{V}$ is a single-context explanation of $\mathcal{N}(x)$ according to Definition 4.1 iff $\vec{V}_{exp}=0$ (that is, all variables in $\vec{V}_{exp}$ have the value 0) is an actual cause of $O=0$ in $(M,\vec{u}_{0})$ according to Definition A.1, under the assumption that a fully masked image has a different classification than $x$.
+Proof.
+First we observe that AC1 simply states that in the current context $\vec{u}_{0}$, the value of all variables in $\vec{V}$ is 0, and the classification of the fully masked image is different from $\mathcal{N}(x)$.
+Assume that EXIM1 holds.
+Then, assigning $\vec{V}_{exp}$ to $1$ results in restoring the original classification $\mathcal{N}(x)$, and hence $O=1$, hence AC2 holds with $\vec{W}=\emptyset$.
+For the other direction, assume that AC2 holds.
+Since the variables in $\vec{V}$ are causally independent, changing $\vec{V}_{exp}$ to $1$ does not change any other input variable, and hence we can take $\vec{W}=\emptyset$.
+Then, AC2 implies EXIM1.
+Finally, AC3 is simply the minimality condition, and is equivalent to EXIM2.
+∎
+We note that the theory of actual causality includes the notion of explanations as well (HP01a), but these are defined over a set of contexts, rather than a single context, so are not suitable for our purposes.
+We can, in general, imagine a setting in which we are searching for a subset of pixels of the input image that is sufficient to result in the original classification, regardless of the values of the rest of the pixels—this would be similar to the setting explored in Anchors (anchors).
+However, such a setting leads to significantly less intuitive results for explaining image classification, as also evident from the results of applying Anchors to images.
+Let us now place our definition of sufficient responsibility into the actual causality landscape. For the ease of presentation, we first restate the definition. See 4.2
+In the literature, the degree of responsibility is defined by Chockler and Halpern (CH04) as the quantification of actual causality, hence its precise definition is tied to the definition of the actual cause.
+The latest version is as follows.
+Definition A.3.
+[Degree of responsibility (Hal19)] For $\vec{X}=\vec{x}$ being an actual cause of $\varphi$ in $(M,\vec{u})$, the degree of responsibility of $\vec{X}=\vec{x}$ for the value of $\varphi$ in $(M,\vec{u})$ is defined as
+$dr(\vec{X}=\vec{x},\varphi,M,\vec{u})=\frac{1}{|\vec{X}|+|\vec{W}|},$
+for the smallest $|\vec{X}|+|\vec{W}|$ that satisfy Definition A.1.
+The following is immediate from the definitions, given that in our case $\vec{W}=\emptyset$.
+Proposition A.4.
+Given a image classifier $\mathcal{N}$, an input $x$, a causal model $M$ derived from them as described in Section 4, and a single-context explanation $\vec{V}_{exp}\subseteq\vec{V}$ of $\mathcal{N}(x)$ according to Definition 4.1, the degree of sufficient responsibility of $\vec{V}_{exp}$ for $\varphi$ in $M$ is the degree of responsibility of $O=0$ in $(M,\vec{u}_{0})$ according to Definition A.3, under the assumption that a fully masked image has a different classification than $x$.
+Indeed, the degree of sufficient responsibility of $\vec{V}_{exp}$ is $\frac{1}{\vec{V}_{exp}}$, which is exactly the degree of responsibility according to Definition A.3 for $(M,\vec{u}_{0})$, given that in our model $M$ there is no dependency between the variables.
+Appendix B The Complexity of Our Definitions
+In order to analyze the complexity of our definition of single-context explanation, we first need to introduce a complexity class.
+Papadimitriou and Yannakakis (PY84) introduced the complexity class $DP$, which consists of all languages $L$ such that there exists a language $L_{1}$ in $NP$ and a language $L_{2}$ in co-$NP$ such that $L=L_{1}\cap L_{2}$.
+Papadimitriou and Yannakakis showed that a number of problems of interest were $DP$ complete.
+For the complexity discussion below, we use the classifier $\mathcal{N}$ as an oracle, not taking its complexity into account.
+Instead, we assume that the value of $O$ is computed in polynomial time using the values of variables in $\vec{V}$.
+Since all variables in our causal models are binary, the computation is simply a Boolean formula.
+Halpern proved that the decision problem of actual causality as per Definition A.1 is $DP$-complete, and the result holds for binary models as well (Hal15).
+The proof, however, relies on causal models having causal dependencies between their variables.
+As we show below, our definition, while still intractable, is in a lower complexity class.
+But first we observe that for singletons, our definition is computable in polynomial time.
+Proposition B.1.
+If $\vec{V}_{exp}$ is a single variable $V$, the complexity of deciding whether it satisfies Definition 4.1 is polynomial in the size of the input.
+Proof.
+The proof is straightforward by observing that EXIM1 is checkable in polynomial time, and EXIM2 holds trivially, as singletons have no subsets. ∎
+We note that the decision problem of actual cause is $NP$-complete for singletons (Hal15), indicating that our setting significantly simplifies the problem.
+The reduction in complexity also holds for the general case, as proved in the following theorem.
+Theorem B.2.
+The decision problem of explanations in Definition 4.1 is co-$NP$-complete.
+Proof.
+For the membership in co-$NP$, we prove that the complementary problem, of showing that a given $\vec{V}_{exp}$ is not an explanation, is in $NP$.
+Given a set $\vec{V}_{exp}$, checking whether it satisfies EXIM1 is polynomial.
+For refuting EXIM2, it suffices to find one witness subset $\vec{V}^{\prime}\subset\vec{V}$ that satisfies EXIM1, hence the complementary problem is indeed in $NP$.
+For the co-NP-hardness, we show a reduction from the classic co-$NP$-complete problem $UNSAT=\{$ unsatisfiable propositional Boolean formulae }\}.
+Given a propositional Boolean formula $\varphi$ over the set of variables $X_{1},\ldots,X_{n}$, we construct a depth-$2$ binary causal model $M_{\varphi}$ as follows.
+The set of endogenous variables $\mathcal{}V$} is the set $X_{1},\ldots,X_{n},O$, where the values of $X_{1},\ldots,X_{n}$ are determined directly by the exogenous variables, and
+$O=(\varphi\vee\ \bigwedge_{i=1}^{n}X_{i})\wedge(\bigwedge_{i=1}^{n}X_{i}\rightarrow\neg{\varphi}).$
+Essentially, the value of $O$ is an exclusive OR of the value of $\varphi$ and the conjunction of all variables of $\varphi$.
+Then, $\varphi\in UNSAT$ iff the set $\vec{X}=\{X_{1},\ldots X_{n}\}$ is a single-context explanation for $O=0$ in $(M,\vec{u}_{0})$.
+First, we claim that $O$ has the value $1$ only in the assignment that sets all variables in $\vec{X}$ to $1$ iff $\varphi$ is unsatisfiable.
+Indeed, if $\varphi$ is unsatisfiable, then its value is 0 under all assignments, and the value of the conjunction is $1$ iff all variables in $\vec{X}$ are set to $1$.
+For the other direction, if the value of $O$ is $1$ under the assignment that sets all variables in $\vec{X}$ to $1$, then $\varphi$ is falsified under this assignment, and since this is the only assignment that results in $O=1$, $\varphi$ is falsified under all other assignments to $\vec{X}$ as well, hence $\varphi$ is unsatisfiable.
+We are now ready to prove the reduction.
+Assume first that $\varphi\in UNSAT$.
+Then, by the previous discussion, $(M,\vec{u}_{0})\models[\vec{X}\leftarrow 1](O=1)$, satisfying EXIM1, and $\vec{X}$ is the only such subset of variables, satisfying EXIM2.
+Therefore, $\vec{X}$ is a single-context explanation to $O$ in $M_{\varphi}$, as required.
+For the other direction, assume that $\vec{X}$ is a single-context explanation to $O$ in $M_{\varphi}$.
+Then, $(M,\vec{u}_{0})\models[\vec{X}\leftarrow 1](O=1)$, and no subset of $\vec{X}$ satisfies this condition, hence by the previous discussion, $\varphi\in UNSAT$, completing the proof.
+∎
+There is, therefore, little hope to find an efficient algorithm for computing exact explanations for image classification, and the answer to RQ2 wrt precise explanations is ‘no’, as computing explanations is intractable.
+While this is not directly relevant to this paper, the following observation is a direct corollary from Lemma A.2 and Theorem B.2.
+Observation B.3.
+In depth-2 binary causal models, the decision problem of actual causality is co-$NP$-complete.
+We now turn to examining the complexity of the degree of sufficient responsibility.
+First, we need to introduce a new complexity hierarchy, specifically, the hierarchy of functional problems.
+This is because the output of the degree of the responsibility computation is a value, and not an accept/reject decision.
+For a complexity class $A$, the class $\mbox{FP}^{{\rm A}[\log{n}]}$ consists of all functions that can be computed by a polynomial-time Turing machine with an oracle for a problem in $A$, which on input $x$ asks a total of $O(\log{|x|})$ queries (cf.
+(Pap84)).
+A function $f(x)$ is $\mbox{FP}^{{\rm A}[\log{n}]}$ hard iff for every function $g(x)$ in $\mbox{FP}^{{\rm A}[\log{n}]}$ there exist polynomially computable functions $R,S:\Sigma^{*}\rightarrow\Sigma^{*}$ (where $\Sigma$ is the common alphabet) such that $g(x)=S(f(R(x)))$.
+A function $f(x)$ is complete in $\mbox{FP}^{{\rm A}[\log{n}]}$ iff it is in $\mbox{FP}^{{\rm A}[\log{n}]}$ and is $\mbox{FP}^{{\rm A}[\log{n}]}$-hard.
+We prove the following result.
+Theorem B.4.
+The problem of computing the degree of sufficient responsibility is $\mbox{FP}^{{\rm NP}[\log{n}]}$-complete.
+Proof.
+The proof is very similar to the proof of complexity of the degree of responsibility in (CH04), so we present only the key points here.
+For the membership in $\mbox{FP}^{{\rm NP}[\log{n}]}$, we observe that the polynomial time algorithm for computing the degree of responsibility can perform a binary search on the size of $\vec{V}_{exp}$, in each step querying an oracle for the existence of a single-context explanation $\vec{V}_{exp}$ of size $k$.
+By  Theorem B.2, the decision problem of the existence of explanation is co-$NP$-complete, and hence can be decided by an $NP$-oracle.
+For hardness in $\mbox{FP}^{{\rm NP}[\log{n}]}$, the reduction is from the $\mbox{FP}^{{\rm NP}[\log{n}]}$-complete function problem $MINSAT$, defined in (CH04) as $MINSAT(\varphi)=k$, where $k$ is the minimal number of $1$’s in a satisfying assignment for $\varphi$.
+The reduction is the same as in (CH04).
+∎
+Theorem B.4 shows that computing the degree of sufficient responsibility in our setting is intractable, hence the need for approximation algorithms.